### v1

# Provably Efficient Reinforcement Learning with Multinomial Logit Function Approximation

Abstract
We study a new class of MDPs that employs multinomial logit function approximation to ensure valid probability distributions over the state space. Despite its benefits, introducing the non-linear function raises significant challenges in both computational and statistical efficiency. The best-known method of Hwang and Oh (2023) has achieved an $\widetilde{\mathcal{O}}(\kappa^{-1}dH^{2}\sqrt{K})$ regret, where $\kappa$ is a problem-dependent quantity, $d$ is the feature space dimension, $H$ is the episode length, and $K$ is the number of episodes. While this result attains the same rate in $K$ as the linear cases, the method requires storing all historical data and suffers from an $\mathcal{O}(K)$ computation cost per episode. Moreover, the quantity $\kappa$ can be exponentially small, leading to a significant gap for the regret compared to the linear cases. In this work, we first address the computational issue by proposing an online algorithm that achieves the same regret with only $\mathcal{O}(1)$ computation cost. Then, we design two algorithms that leverage local information to enhance statistical efficiency. They not only maintain an $\mathcal{O}(1)$ computation cost per episode but achieve improved regrets of $\widetilde{\mathcal{O}}(\kappa^{-1/2}dH^{2}\sqrt{K})$ and $\widetilde{\mathcal{O}}(dH^{2}\sqrt{K}+\kappa^{-1}d^{2}H^{2})$ respectively. Finally, we establish a lower bound, justifying the optimality of our results in $d$ and $K$. To the best of our knowledge, this is the first work that achieves almost the same computational and statistical efficiency as linear function approximation while employing non-linear function approximation for reinforcement learning.
1 Introduction
Reinforcement Learning (RL) with function approximation has achieved great success in many applications with large state and action space, such as games (Mnih et al., 2013), robotics (Kober et al., 2013), large language models (Ouyang et al., 2022), etc. Therefore, establishing the theoretical foundation for RL with function approximation is of great importance. Recently, there have been many efforts devoted to understanding the linear function approximation due to its significance and simplicity, yielding fruitful results (Yang and Wang, 2019; Jin et al., 2020; Ayoub et al., 2020).
While these works make important steps toward understanding reinforcement learning with function approximation, there are still some challenges to be solved. In RL with linear function approximation, transitions are presumed to be linear in specified feature mappings, such as $\mathbb{P}(s^{\prime}|s,a)=\phi(s^{\prime}|s,a)^{\top}\theta^{*}$ for linear mixture MDPs and $\mathbb{P}(s^{\prime}|s,a)=\phi(s,a)^{\top}\mu^{*}(s^{\prime})$ for linear MDPs. Here $\mathbb{P}(s^{\prime}|s,a)$ is the probability from state $s$ to state $s^{\prime}$ taking action $a$, $\phi(s^{\prime}|s,a)$ and $\phi(s,a)$ are feature mappings, $\theta^{*}$ and $\mu^{*}(s^{\prime})$ are unknown parameters. However, a transition function is a probability distribution over states, i.e., the magnitude falls within the range of $[0,1]$, and the aggregate probability equals $1$. Thus, for certain feature mappings, linear transitions may not yield a valid probability distribution with arbitrary $\theta$ and $\mu$. In general, an algorithm designed for linear MDPs could break down entirely if the underlying MDP is not linear as noted by Jin et al. (2020). While there are some works exploring generalized linear function approximation (Wang et al., 2021) and general function approximation (Russo and Roy, 2013; Du et al., 2021; Foster et al., 2021), they make assumptions over value functions rather than transitions, hence do not tackle this challenge.
Table 1: Comparison between previous work and ours in terms of the regret bound, computation cost and storage cost. Here $\kappa$ and $\kappa^{*}$ are exponentially small problem-dependent quantities, $d$ is the feature dimension, $H$ is the episode length and $K$ is the number of episodes. The computational cost and storage cost per episode indicate the dependence on episode count $k$.
Reference Regret Comp. Stor. MDP model
Hwang and Oh (2023) $\widetilde{\mathcal{O}}(\kappa^{-1}dH^{2}\sqrt{K})$ $\mathcal{O}(k)$ $\mathcal{O}(k)$ homogeneous
UCRL-MNL-OL (Theorem 1) $\widetilde{\mathcal{O}}(\kappa^{-1}dH^{2}\sqrt{K})$ $\mathcal{O}(1)$ $\mathcal{O}(1)$ inhomogeneous
UCRL-MNL-LL (Theorem 2) $\widetilde{\mathcal{O}}(\kappa^{-1/2}dH^{2}\sqrt{K})$ $\mathcal{O}(1)$ $\mathcal{O}(1)$ inhomogeneous
UCRL-MNL-LL+ (Theorem 3) $\widetilde{\mathcal{O}}(dH^{2}\sqrt{K}+\kappa^{-1}d^{2}H^{2})$ $\mathcal{O}(1)$ $\mathcal{O}(1)$ inhomogeneous
Lower Bound  (Corollary 1) $\Omega(dH\sqrt{K\kappa^{*}})$ / / infinite action space
Towards addressing the limitation of linear function approximation, a new class of MDPs that utilizes multinomial logit (MNL) function approximation has been proposed by Hwang and Oh (2023) recently. Despite its benefits, introducing non-linear functions raises significant challenges in both computational and statistical efficiency. Specifically, the best-known method of Hwang and Oh (2023) has achieved an $\widetilde{\mathcal{O}}(\kappa^{-1}dH^{2}\sqrt{K})$ regret. While matching the optimal regret of $\widetilde{\mathcal{O}}(d\sqrt{H^{3}K})$ as linear function approximation in terms of $K$ (Zhou et al., 2021; He et al., 2023), their method requires storing all historical data and the computational cost per episode grows linearly with the episode count, which is inefficient. Moreover, the quantity $\kappa$ can be exponentially small, leading to a significant gap for the regret compared to the linear cases. Thus, a natural question arises:
Can we achieve (almost) the same computational and statistical efficiency as linear function approximation while employing non-linear function approximation for RL?
In this paper, we answer this question affirmatively by making significant advancements in both computational and statistical efficiency for RL with MNL function approximation, almost matching the results of linear function approximation. Our contributions can be summarized as follows:
  • We first propose the UCRL-MNL-OL algorithm based on a variant of Online Newton Step, attaining an $\widetilde{\mathcal{O}}(\kappa^{-1}dH^{2}\sqrt{K})$ regret with $\mathcal{O}(1)$ computational and storage costs per episode. It matches the best-known regret by Hwang and Oh (2023), yet achieves the same computational and storage efficiency as linear function approximation (Jin et al., 2020; Zhou et al., 2021).
  • As the quantity $\kappa$ can be exponentially small, we propose two algorithms UCRL-MNL-LL and UCRL-MNL-LL+, which leverage local information to improve the statistical efficiency. They not only maintain $\mathcal{O}(1)$ computational and storage costs but also achieve improved regrets of $\widetilde{\mathcal{O}}(\kappa^{-1/2}dH^{2}\sqrt{K})$ and $\widetilde{\mathcal{O}}(dH^{2}\sqrt{K}+d^{2}H^{2}\kappa^{-1})$, respectively. The latter result moves the $\kappa$ dependence to the lower-order term which is independent of $K$ and the higher-order term matches the optimal regret $\widetilde{\mathcal{O}}(d\sqrt{H^{3}K})$ for linear mixture MDPs (Zhou et al., 2021) and linear MDPs (He et al., 2023) in both $d$ and $K$, differing only by an $\mathcal{O}(H^{1/2})$ factor.
  • We establish the first lower bound for MDPs with MNL function approximation by employing a reduction to the logistic bandits. We prove a lower bound of $\Omega(dH\sqrt{K\kappa^{*}})$ for the infinite action space setting. Though this does not constitute a strict lower bound for the finite actions case studied in this work, it justifies that our results may be optimal in $d$ and $K$.
Table 1 shows comparisons between previous works and ours. From a technical perspective, inspired by recent advances in multinomial logistic bandits (Zhang and Sugiyama, 2023), we estimate the parameter by running an online algorithm instead of performing maximum likelihood estimation over all historical data. Differently, we need to design value functions carefully that utilize local information to address the unique challenges of MDPs, which is highly non-trivial.
Organization.
We first introduce the related work in Section 2 and present the setup in Section 3. Then, we introduce an algorithm with constant computational and storage costs in Section 4. Next, we present two improved algorithms in Section 5. Finally, we establish the lower bound in Section 6. Section 7 concludes the paper. Due to space constraints, we defer all proofs to the appendixes.
Notations.
We denote by $[n]$ the set $\{1,\ldots,n\}$ and define $[x]_{[a,b]}=\min(\max(x,a),b)$. For $X,Y\in\mathbb{R}^{d\times d}$, $X\succeq Y$ means $X-Y$ is positive semi-definite. For a vector $\mathbf{x}\in\mathbb{R}^{d}$ and positive semi-definite matrix $A\in\mathbb{R}^{d\times d}$, denote $\lVert\mathbf{x}\rVert_{A}=\sqrt{\mathbf{x}^{\top}A\mathbf{x}}$. The $\widetilde{\mathcal{O}}(\cdot)$ hides all polylogarithmic factors.
2 Related Work
In this section, we review related works, both in terms of settings and techniques.
RL with Generalized Linear Function Approximation.
There are recent efforts devoted to investigating function approximation beyond the linear models. Wang et al. (2021) investigated RL with generalized linear function approximation. Notably, unlike our approach which models transitions using a generalized linear model, they apply this approximation directly to the value function. Another line of works (Chowdhury et al., 2021; Li et al., 2022; Ouhamma et al., 2023) has studied RL with exponential function approximation and also aimed to ensure that transitions constitute valid probability distributions. The MDP model can be viewed as an extension of bilinear MDPs in their work while our setting extends linear mixture MDPs. These studies are complementary to ours and not directly comparable. Moreover, these works also enter the computational and statistical challenges arising from non-linear function approximation that remain to be addressed. The most relevant work to ours is the recent work by Hwang and Oh (2023), which firstly explored a similar setting to ours, where the transition is modeled using a multinomial logit model. We significantly improve their results by providing computationally and statistically efficient algorithms.
RL with General Function Approximation.
There have also been some works studies reinforcement learning with general function approximation. Russo and Roy (2013) and Osband and Roy (2014) initiated the study on the minimal structural assumptions that render sample-efficient learning by proposing a structural condition called Eluder dimension. Recently, several works have investigated different conditions for sample-efficient interactive learning, such as Bellman Eluder (BE) dimension (Jin et al., 2021), Bilinear classes (Du et al., 2021), Decision-Estimation Coefficient (DEC) (Foster et al., 2021), and Admissible Bellman Characterization (ABC) (Chen et al., 2023). A notable difference is that they impose assumptions on the value functions while we study function approximation on the transitions. Moreover, the goal of these works is to study the conditions for sample-efficient reinforcement learning, but not focus on the computational efficiency.
Multinomial Logit (MNL) Contextual Bandits.
As our work shares similar techniques with MNL contextual bandits to some extent, we briefly review works on this topic. The pioneering study by Cheung and Simchi-Levi (2017) introduced an algorithm that achieves a Bayesian regret of $\widetilde{\mathcal{O}}(\kappa^{-1}d\sqrt{K})$, where $K$ denotes the number of rounds. This result was further enhanced by subsequent studies (Oh and Iyengar, 2019, 2021; Agrawal et al., 2023). The recent work by Périvier and Goyal (2022) has significantly improved the dependence on $\kappa$, obtaining a regret of $\widetilde{\mathcal{O}}(d\sqrt{\kappa K}+\kappa^{-1})$ in the uniform revenue setting. Most of the above methods require storing all historical data and are computationally inefficient. A concurrent work with ours by Lee and Oh (2024) proposed an algorithm with constant computational and storage costs inspired by recent advancements in multinomial logistic bandits (Zhang and Sugiyama, 2023) and achieved optimal regret of $\widetilde{\mathcal{O}}(d\sqrt{\kappa K})$ and $\widetilde{\mathcal{O}}(d\sqrt{K})$ under uniform and nonuniform rewards respectively. However, although the underlying models of MNL contextual bandits and MNL mixture MDPs are similar, the challenges they pose differ significantly and the techniques developed for MNL contextual bandits can not be directly applied to the MNL mixture MDPs. Specifically, in MNL bandits, the objective is to select a series of assortments with varying sizes that maximize expected revenue, whereas in MNL mixture MDPs, the goal is to maximize cumulative rewards Therefore, while we utilize similar parameter estimation techniques as those in MNL contextual bandits, we need to construct value functions to address the unique challenges of MDPs, which is highly non-trivial and needs a careful design.
3 Problem Setup
In this section, we present the problem setup of RL with multinomial logit function approximation.
Inhomogeneous, Episodic MDPs.
An inhomogeneous, episodic MDP instance can be denoted by a tuple $M=(\mathcal{S},\mathcal{A},H,\{\mathbb{P}_{h}\}_{h=1}^{H},\{r_{h}\}_{h=1}^{H})$, where $\mathcal{S}$ is the state space, $\mathcal{A}$ is the action space, $H$ is the length of each episode, $\mathbb{P}_{h}:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\to[0,1]$ is the transition kernel with $\mathbb{P}_{h}(s^{\prime}\>|\>s,a)$ is being the probability of transferring to state $s^{\prime}$ from state $s$ and taking action $a$ at stage $h$, $r_{h}:\mathcal{S}\times\mathcal{A}\to[0,1]$ is the deterministic reward function. A policy $\pi=\{\pi_{h}\}_{h=1}^{H}$ is a collection of mapping $\pi_{h}$, where each $\pi_{h}:\mathcal{S}\to\Delta(\mathcal{A})$ is a function maps a state $s$ to distributions over $\mathcal{A}$ at stage $h$. For any policy $\pi$ and $(s,a)\in\mathcal{S}\times\mathcal{A}$, we define the action-value function $Q_{h}^{\pi}$ and value function $V_{h}^{\pi}$ as follows:
$Q_{h}^{\pi}(s,a)=\mathbb{E}\left[\sum_{h^{\prime}=h}^{H}r_{h^{\prime}}\left(s_{h^{\prime}},a_{h^{\prime}}\right)\Big{|}s_{h}=s,a_{h}=a\right],\quad V_{h}^{\pi}(s)=\mathbb{E}_{a\sim\pi_{h}(\cdot\mid s)}\left[Q_{h}^{\pi}(s,a)\right],$
where the expectation of $Q_{h}^{\pi}$ is taken over the randomness of the transition $\mathbb{P}$ and policy $\pi$. The optimal value function $V_{h}^{*}$ and action-value function $Q_{h}^{*}$ given by $V_{h}^{*}(s)=\sup_{\pi}V_{h}^{\pi}(s)$ and $Q_{h}^{*}(s,a)=\sup_{\pi}Q_{h}^{\pi}(s,a)$. For any function $V:\mathcal{S}\to\mathbb{R}$, we define $[\mathbb{P}_{h}V](s,a)=\mathbb{E}_{s^{\prime}\sim\mathbb{P}_{h}(\cdot\mid s,a)}V(s^{\prime})$. The Bellman equation for policy $\pi$ and the Bellman optimality equation are given by
$Q_{h}^{\pi}(s,a)=r_{h}(s,a)+\left[\mathbb{P}_{h}V_{k,h+1}^{\pi}\right](s,a),\quad Q_{h}^{*}(s,a)=r_{h}(s,a)+\left[\mathbb{P}_{h}V_{k,h+1}^{*}\right](s,a).$
Learning Protocol.
In the online MDP setting, the learner interacts with the environment without the knowledge of the transition $\{\mathbb{P}_{h}\}_{h=1}^{H}$. For the sake of simplicity, we assume the reward $\{r_{h}\}_{h=1}^{H}$ is known. The interaction proceeds in $K$ episodes. At the beginning of each episode $k$, the learner chooses a policy $\pi_{k}=\{\pi_{k,h}\}_{h=1}^{H}$. At each stage $h\in[H]$, starting from the initial state $s_{k,1}$, the learner observes the state $s_{k,h}$, chooses an action $a_{k,h}$ sampled from $\pi_{k,h}(\cdot\>|\>s_{k,h})$, obtains reward $r_{h}(s_{k,h},a_{k,h})$ and transits to the next state $s_{k,h+1}\sim\mathbb{P}_{h}(\cdot\>|\>s_{k,h},a_{k,h})$ for $h\in[H]$. The episode ends when $s_{H+1}$ is reached. The goal of the learner is to minimize regret, defined as the difference between the cumulative reward of the optimal policy and the learner’s policy, namely,
$\operatorname*{Reg}(K)=\sum_{k=1}^{K}V_{1}^{*}(s_{k,1})-\sum_{k=1}^{K}V_{1}^{\pi_{k}}(s_{k,1}).$
Multinomial Logit (MNL) Mixture MDPs.
Although significant advances have been achieved for MDPs with linear function approximation, Hwang and Oh (2023) show that there exists a set of features such that no linear transition model (including bilinear and low-rank MDPs) can induce a valid probability distribution over the state space, which limits the expressiveness of function approximation. To overcome this limitation, they propose a new class of MDPs with multinomial logit function approximation. We introduce the formal definition of MNL mixture MDPs below.
Definition 1 (Reachable States).
For any $(h,s,a)\in[H]\times\mathcal{S}\times\mathcal{A}$, we define the “reachable states” as the set of states that can be reached from state $s$ taking action $a$ at stage $h$ within a single transition, i.e., $\mathcal{S}_{h,s,a}\triangleq\{s^{\prime}\in\mathcal{S}\mid\mathbb{P}_{h}(s^{\prime}\>|\>s,a)>0\}$. Also, we define $S_{h,s,a}\triangleq|\mathcal{S}_{h,s,a}|$ and further define $U\triangleq\max_{(h,s,a)}S_{h,s,a}$ the maximum number of reachable states.
Definition 2 (MNL Mixture MDP).
An MDP instance $M=(\mathcal{S},\mathcal{A},H,\{\mathbb{P}_{h}\}_{h=1}^{H},\{r_{h}\}_{h=1}^{H})$ is called an inhomogeneous, episodic $B$-bounded MNL mixture MDP if there exist a known feature mapping $\phi(s^{\prime}\>|\>s,a):\mathcal{S}\times\mathcal{A}\times\mathcal{S}\to\mathbb{R}^{d}$ with $\lVert\phi(s^{\prime}\>|\>s,a)\rVert_{2}\leq 1$ and unknown vectors $\{\theta_{h}^{*}\}_{h=1}^{H}\in\Theta$ with $\Theta=\{\theta\in\mathbb{R}^{d},\lVert\theta\rVert_{2}\leq B\}$, such that for all $(s,a,h)\in\mathcal{S}\times\mathcal{A}\times[H]$ and $s^{\prime}\in\mathcal{S}_{h,s,a}$, it holds that
$\mathbb{P}_{h}(s^{\prime}\>|\>s,a)=\frac{\exp(\phi(s^{\prime}\>|\>s,a)^{\top}\theta_{h}^{*})}{\sum_{\widetilde{s}\in\mathcal{S}_{h,s,a}}\exp(\phi(\widetilde{s}\>|\>s,a)^{\top}\theta_{h}^{*})}.$
For any $\theta\in\mathbb{R}^{d}$, we define the induced transition as $p_{s,a}^{s^{\prime}}(\theta)=\frac{\exp(\phi(s^{\prime}\>|\>s,a)^{\top}\theta)}{\sum_{\widetilde{s}\in\mathcal{S}_{s,a}}\exp(\phi(\widetilde{s}\>|\>s,a)^{\top}\theta)}$. Next, we introduce a key assumption for this problem as in Hwang and Oh (2023).
Assumption 1.
There exists $10<\kappa\leq\kappa^{*}<$ 1 such that for all $(s,a,h)\in\mathcal{S}\times\mathcal{A}\times[H]$ and $s^{\prime},s^{\prime\prime}\in\mathcal{S}_{h,s,a}$, it holds that $\inf_{\theta\in\Theta}p_{s,a}^{s^{\prime}}(\theta)p_{s,a}^{s^{\prime\prime}}(\theta)\geq\kappa$ and $p_{s,a}^{s^{\prime}}(\theta^{*})p_{s,a}^{s^{\prime\prime}}(\theta^{*})>\kappa^{*}$.
Assumption 1 is equivalent to the standard assumption in generalized linear model literature (Filippi et al., 2010; Wang et al., 2021) and logistic bandits literature (Faury et al., 2020; Abeille et al., 2021; Zhang and Sugiyama, 2023) to guarantee the Hessian matrix is non-singular.
4 Efficient Algorithm with Online Update
In this part, we design an algorithm that achieves $\mathcal{O}(1)$ computational and storage costs.
Efficiency Concern in Previous Methods.
While Hwang and Oh (2023) made the first step and proposed the UCRL-MNL algorithm with $\widetilde{\mathcal{O}}(\kappa^{-1}dH^{2}\sqrt{K})$ regret, the computational and storage complexity of their algorithm scales linearly with the episode count, which is expensive. Specifically, for the trajectory $\{(s_{k,h},a_{k,h})\}_{h=1}^{H}$ at episode $k$, we define the variable: $y_{k,h}\in\{0,1\}^{N_{k,h}}$ where $y_{k,h}^{s^{\prime}}=\operatorname{\mathds{1}}{(s_{k,h+1}=s^{\prime})}$ for $s^{\prime}\in\mathcal{S}_{k,h}\triangleq\mathcal{S}_{s_{k,h},a_{k,h}}$ and $N_{k,h}=\lvert\mathcal{S}_{k,h}\rvert$. Furthermore, we denote by $p_{k,h}^{s^{\prime}}(\theta)=p_{s_{k,h},a_{k,h}}^{s^{\prime}}(\theta)$. Then $y_{k,h}$ is a sample from the following multinomial distribution:
$y_{k,h}\sim\textnormal{multinomial}(1,[p_{k,h}^{s_{1}}(\theta^{*}),\ldots,p_{k,h}^{s_{N_{k,h}}}(\theta^{*})]).$
where the parameter $1$ indicates that $y_{k,h}$ is a single-trial sample. We define the noise $\epsilon_{k,h}^{s^{\prime}}=y_{k,h}^{s^{\prime}}-p_{k,h}^{s^{\prime}}(\theta_{h}^{*})$. It is clear that $\epsilon_{k,h}\in[-1,1]^{N_{k,h}}$, $\mathbb{E}[\epsilon_{k,h}]=\mathbf{0}$ and $\sum_{s^{\prime}\in\mathcal{S}_{k,h}}\epsilon_{i,h}^{s^{\prime}}=0$.
The UCRL-MNL algorithm estimates the unknown transition parameter $\theta_{h}^{*}$ using the regularized maximum likelihood estimation (MLE). Specifically, based on $y_{k,h}$, the estimator $\theta_{k,h}$ is defined as the solution to the following optimization problem:
$\theta_{k,h}=\operatorname*{arg\,min}_{\theta\in\mathbb{R}^{d}}\sum_{i=1}^{k-1}\sum_{s^{\prime}\in\mathcal{S}_{i,h}}-y_{i,h}^{s^{\prime}}\log p_{i,h}^{s^{\prime}}(\theta)+\frac{\lambda}{2}\lVert\theta\rVert_{2}^{2}.$ (1)
In episode $k$, the estimator $\theta_{k,h}$ in (1) is computed using all samples collected in previous episodes, resulting in the computational and storage cost per episode that scales linearly with the episode count, which is expensive in practice.
To address this issue, we introduce a novel algorithm named UCRL-MNL with Online Learning (UCRL-MNL-OL), which achieves both the same regret with constant computational and storage cost. At a high level, our algorithm can be divided into two phases: (i) efficient online estimation, and (ii) efficient optimistic value function construction. We introduce the details in the following.
Efficient Online Estimation.
Instead of using all historical data, our algorithm updates the estimator in an efficient online manner. This is achieved by efficiently exploiting particular properties of the MNL model. The key observation is that the negative log-likelihood function is exponentially concave over the bounded domain, which motivates us to apply a variant of online Newton step (Hazan et al., 2007), in particular, inspired by the works (Hazan et al., 2014; Zhang et al., 2016; Oh and Iyengar, 2021), which focus on logistic bandits and multinomial logit contextual bandits.
First, we define pe-episode loss function $f_{k,h}(\theta)$, gradient $g_{k,h}(\theta)$ and Hessian matrix $H_{k,h}(\theta)$ as
$f_{k,h}(\theta)$ $=-\sum_{s^{\prime}\in\mathcal{S}_{k,h}}y_{k,h}^{s^{\prime}}\log p_{k,h}^{s^{\prime}}(\theta),\quad g_{k,h}(\theta)=\sum_{s^{\prime}\in\mathcal{S}_{k,h}}(p_{k,h}^{s^{\prime}}(\theta)-y_{k,h}^{s^{\prime}})\phi_{k,h}^{s^{\prime}},$
$H_{k,h}(\theta)$ $=\sum_{s^{\prime}\in\mathcal{S}_{k,h}}p_{k,h}^{s^{\prime}}(\theta)\phi_{k,h}^{s^{\prime}}(\phi_{k,h}^{s^{\prime}})^{\top}-\sum_{s^{\prime}\in\mathcal{S}_{k,h}}\sum_{s^{\prime\prime}\in\mathcal{S}_{k,h}}p_{k,h}^{s^{\prime}}(\theta)p_{k,h}^{s^{\prime\prime}}(\theta)\phi_{k,h}^{s^{\prime}}(\phi_{k,h}^{s^{\prime\prime}})^{\top}.$ (2)
We introduce the following key property of the loss function $f_{k,h}(\theta)$ by a second-order Taylor expansion, whose proof can be found in Lemma 5 in Appendix A.
Property 1.
(i). For any $\theta_{1},\theta_{2}\in\mathbb{R}^{d}$, $\exists~{}\bar{\theta}=\nu\theta_{1}+(1-\nu)\theta_{2}$ such that
$f_{k,h}(\theta_{2})=f_{k,h}(\theta_{1})+\langle g_{k,h}(\theta_{1}),\theta_{2}-\theta_{1}\rangle+\frac{1}{2}\lVert\theta_{2}-\theta_{1}\rVert_{H_{k,h}(\bar{\theta})}^{2}.$
(ii). For any $\theta\in\Theta$, it holds that
$H_{k,h}(\theta)\succeq\kappa\sum_{s^{\prime}\in\mathcal{S}_{k,h}}\phi_{k,h}^{s^{\prime}}(\phi_{k,h}^{s^{\prime}})^{\top}\triangleq\kappa W_{k,h},\forall\theta\in\Theta.$ (3)
This implies that $f_{k,h}(\theta)$ is exponentially concave over the bounded domain $\Theta$, which enables us to apply the online Newton step to construct an efficient estimator. Specifically, we construct the estimator $\widehat{\theta}_{k+1,h}$ by solving the problem:
$\widehat{\theta}_{k+1,h}=\operatorname*{arg\,min}_{\theta\in\Theta}\langle g_{k,h}(\widehat{\theta}_{k,h}),\theta-\widehat{\theta}_{k,h}\rangle+\frac{1}{2}\lVert\theta-\widehat{\theta}_{k,h}\rVert_{\widehat{\Sigma}_{k+1,h}}^{2},$ (4)
where $\widehat{\Sigma}_{k+1,h}=\widehat{\Sigma}_{k,h}+\frac{\kappa}{2}\sum_{s^{\prime}\in\mathcal{S}_{k,h}}\phi_{k,h}^{s^{\prime}}(\phi_{k,h}^{s^{\prime}})^{\top}$ is the feature covariance matrix. This estimator can be computed efficiently in an online manner without the need to store all historical data. Then, we construct the confidence set by the following lemma.
Lemma 1.
For any $\delta\in(0,1)$, define the confidence set as
$\widehat{\mathcal{C}}_{k,h}=\bigg{\{}\theta\in\Theta\mid\lVert\theta-\widehat{\theta}_{k,h}\rVert_{\widehat{\Sigma}_{k,h}}\leq\sqrt{\lambda}B+\sqrt{\frac{8}{\kappa}d\log\left(1+\frac{kUH}{d\lambda\delta}\right)}\triangleq\widehat{\beta}_{k}\bigg{\}}.$
Then, we have $\Pr[\theta_{h}^{*}\in\widehat{\mathcal{C}}_{k,h}]\geq 1-\delta,\forall k\in[K],h\in[H]$.
Remark 1.
Notably, a similar confidence set is achieved by Hwang and Oh (2023) by using the MLE defined in (1). In contrast, we obtain the same results by using an online estimator which only suffers $\mathcal{O}(1)$ computation and storage cost per round. Besides, we identify a technical issue in their analysis. Specifically, they bound the confidence set using the self-normalized concentration for vector-valued martingales in Lemma 15. However, it is crucial to recognize that the noise is not independent and satisfies $\sum_{s^{\prime}\in\mathcal{S}_{i,h}}\epsilon_{i,h}^{s^{\prime}}=0$ since the learner visits each stage $h$ exactly once per episode. Thus, the noise $\epsilon_{i,h}^{s^{\prime}}$ becomes deterministic and non-zero given the remaining noise $\epsilon_{i,h}^{s^{\prime\prime}}$ for $s^{\prime\prime}\neq s^{\prime}$. This contravenes the zero-mean sub-Gaussian condition in Lemma 15. We observe similar oversights also appear in the works on multinomial logit contextual bandits (Oh and Iyengar, 2019, 2021; Agrawal et al., 2023). To the best of our knowledge, this issue has not been explicitly identified in previous studies. We note that this issue can be resolved with only slight modifications in constant factors by a new self-normalized concentration with dependent noises in Lemma 1 of Li et al. (2024), a simplified version of Theorem C.6 in Périvier and Goyal (2022).
Algorithm 1 UCRL-MNL-OL
0:  Regularization parameter $\lambda$, confidence width $\widehat{\beta}_{k}$, confidence parameter $\delta$.
1:  Initialization:  $\widehat{\Sigma}_{1,h}=\lambda I,\widehat{\theta}_{1,h}=\mathbf{0}$ for all $h\in[H]$.
2:  for $k=1,\ldots,K$ do
3:     Compute $\widehat{Q}_{k,h}(\cdot,\cdot)$ in a backward way as in (6).
4:     for $h=1,\ldots,H$ do
5:        Observe current state $s_{k,h}$, select action $a_{k,h}=\operatorname*{arg\,max}_{a\in\mathcal{A}}\widehat{Q}_{k,h}(s_{k,h},a)$.
6:        Set $\widehat{\Sigma}_{k+1,h}=\widehat{\Sigma}_{k,h}+\frac{\kappa}{2}\sum_{s^{\prime}\in\mathcal{S}_{k,h}}\phi_{k,h}^{s^{\prime}}(\phi_{k,h}^{s^{\prime}})^{\top}$.
7:        Compute $\widehat{\theta}_{k+1,h}=\operatorname*{arg\,min}_{\theta\in\Theta}\langle g_{k,h}(\widehat{\theta}_{k,h}),\theta-\widehat{\theta}_{k,h}\rangle+\frac{1}{2}\lVert\theta-\widehat{\theta}_{k,h}\rVert_{\widehat{\Sigma}_{k+1,h}}^{2}$ as (4).
8:     end for
9:  end for
Efficient Optimistic Value Function Construction.
Given the confidence set $\widehat{\mathcal{C}}_{k,h}$, the most direct way is to adopt the principle of “optimism in the face of uncertainty” and construct the optimistic value function as the maximum expected reward over the confidence set. Specifically, define the value function $\widehat{Q}_{k,h}(s,a)$ and $\widehat{V}_{k,h}(s)$ as
$\widehat{Q}_{k,h}(s,a)=\bigg{[}r_{h}(s,a)+\max_{\theta\in\widehat{\mathcal{C}}_{k,h}}\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\theta)\widehat{V}_{k,h+1}(s^{\prime})\bigg{]}_{[0,H]},\widehat{V}_{k,h}(s)=\max_{a\in\mathcal{A}}\widehat{Q}_{k,h}(s,a).$ (5)
However, this construction is not efficient as it requires solving a maximization problem over the confidence set. Actually, the algorithm only requires the estimated action-value function to be an upper bound for the true value functions. Thus, we can use a closed-form confidence bound instead of computing the maximal value over the confidence set. To this end, we present the following lemma, which enables us to construct the optimistic value function efficiently.
Lemma 2.
Suppose Lemma 1 holds. For any function $V:\mathcal{S}\to[0,H]$ and $(h,s,a)\in[H]\times\mathcal{S}\times\mathcal{A}$, with probability at least $1-\delta$, it holds that
$\bigg{|}{\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\widehat{\theta}_{k,h})V(s^{\prime})-\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\theta_{h}^{*})V(s^{\prime})}\bigg{|}\leq H\widehat{\beta}_{k}\max_{s^{\prime}\in\mathcal{S}_{h,s,a}}\lVert\phi_{s,a}^{s^{\prime}}\rVert_{\widehat{\Sigma}_{k,h}^{-1}}.$
Based on this lemma, we can replace the maximization problem in (5) with closed-form confidence bound and construct the optimistic value function $\widehat{Q}_{k,h}(s,a)$ as
$\widehat{Q}_{k,h}(s,a)=\bigg{[}r_{h}(s,a)+\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\widehat{\theta}_{k,h})\widehat{V}_{k,h+1}(s^{\prime})+H\widehat{\beta}_{k}\max_{s^{\prime}\in\mathcal{S}_{h,s,a}}\|\phi_{s,a}^{s^{\prime}}\|_{\widehat{\Sigma}_{k,h}^{-1}}\bigg{]}_{[0,H]}.$ (6)
At state $s_{k,h}$, we chooses $a_{k,h}=\operatorname*{arg\,max}_{a\in\mathcal{A}}\widehat{Q}_{k,h}(s_{k,h},a)$. Algorithm 1 presents the detailed procedure. We show that our algorithm enjoys the following regret guarantee.
Theorem 1.
Suppose Assumption 1 holds. For any $\delta\in(0,1)$, set $\widehat{\beta}_{k}$ as in Lemma 1 and $\lambda=1$, with probability at least $1-\delta$, Algorithm 1 ensures the following guarantee
${\operatorname*{Reg}}(K)\leq\widetilde{\mathcal{O}}(\kappa^{-1}dH^{2}\sqrt{K}).$
Remark 2.
Our algorithm attains the same regret of Hwang and Oh (2023), yet for the more general inhomogeneous MDPs. More importantly, our algorithm only requires $\mathcal{O}(1)$ computational and storage costs, matching the computational efficiency of MDPs with linear function approximation (Ayoub et al., 2020; Jin et al., 2020; Zhou et al., 2021; He et al., 2023).
5 Improved Algorithm by Local Learning
While the UCRL-MNL-OL algorithm enjoys favorable computational efficiency, it suffers from a dependence on $\kappa^{-1}$ in the regret. By Assumption 1, the quantity $\kappa$ may be exponentially small, exhibiting an exponential dependence on the radius of the parameter set and linear dependence on the number of reachable states, i.e., $\kappa^{-1}=\mathcal{O}((U\exp(B))^{2})$. Thus, there is a significant gap between MNL function approximation and linear function approximation. To bridge this gap, we propose new algorithms that leverage local information to enhance the dependence on $\kappa$ significantly.
We first analyze where the dependence on $\kappa$ comes from. By the analysis of Theorem 1, we can see that the regret of the algorithm can be upper-bounded as follows,
$\operatorname*{Reg}(K)\leq 2H\widehat{\beta}_{K}\sum_{k=1}^{K}\sum_{h=1}^{H}\max_{s^{\prime}\in\mathcal{S}_{k,h}}\lVert\phi_{k,h}^{s^{\prime}}\rVert_{\widehat{\Sigma}_{k,h}^{-1}}+\sum_{k=1}^{K}\sum_{h=1}^{H}\mathcal{M}_{k,h}.$ (7)
where the first term is the overestimate of the value function and the second term is the martingale difference sequence due to the stochastic transition. The second term can be upper-bounded by the Azuma-Hoeffding inequality which is independent of $\kappa$. Thus, the $\kappa$-dependence comes from the first term. By the construction of the confidence set in Lemma 1, the dependence on $\widehat{\beta}_{k}$ in the first term leads to a $\kappa^{-1/2}$ dependence. By the configuration of the covariance matrix $\widehat{\Sigma}_{k,h}$ in (4), the dependence on $\widehat{\Sigma}_{k,h}$ also leads to a $\kappa^{-1/2}$ dependence. Thus, we need to design a local learning algorithm that can adaptively use local information to enhance the dependence on these two terms.
5.1 Improved Online Estimation
In Section 4, we show the Hessian matrix $H_{k,h}$ is lower bounded by a positive definite matrix $\kappa\sum_{s^{\prime}\in\mathcal{S}_{k,h}}\phi_{k,h}^{s^{\prime}}(\phi_{k,h}^{s^{\prime}})^{\top}$ uniformly over the parameter domain in (3). This lower bound is not tight as the Hessian matrix can be much larger at certain regions rather than the global minimum. This motivates us to design a local learning algorithm that can adaptively use local information.
Inspired by recent advances in multinomial logistic bandit (Zhang and Sugiyama, 2023) and multinomial logit contextual bandit (Lee and Oh, 2024), we run an online algorithm to estimate the parameter $\theta_{h}^{*}$ similar to (4). Differently, we use the local Hessian matrix ~$\widetilde{\mathcal{H}}_{k,h}$ to update the estimation instead of the global lower bound. Specifically, we estimate parameter $\widetilde{\theta}_{k,h}$ as follows:
$\widetilde{\theta}_{k+1,h}=\operatorname*{arg\,min}_{\theta\in\Theta}\langle g_{k,h}(\widetilde{\theta}_{k,h}),\theta-\widetilde{\theta}_{k,h}\rangle+\frac{1}{2\eta}\lVert\theta-\widetilde{\theta}_{k,h}\rVert_{\widetilde{\mathcal{H}}_{k,h}}.$ (8)
where $\eta$ is the step size, $\mathcal{H}_{k,h}=\sum_{i=1}^{k-1}H_{i,h}(\widetilde{\theta}_{i+1,h})+\lambda I_{d}$ and ~$\widetilde{\mathcal{H}}_{k,h}=\mathcal{H}_{k,h}+\eta H_{k,h}(\widetilde{\theta}_{k,h})$. Note both $\mathcal{H}_{k,h}$ and ~$\widetilde{\mathcal{H}}_{k,h}$ can be updated incrementally and the optimization problem in (8) can be solved efficiently. This step can be solved by the following two-step update.
$\theta^{\prime}_{k+1,h}=\theta_{k,h}-\eta\mathcal{H}_{k,h}^{-1}g_{k,h}(\theta_{k,h}),\quad\widetilde{\theta}_{k+1,h}=\operatorname*{arg\,min}_{\theta\in\Theta}\lVert\theta-\theta^{\prime}_{k+1,h}\rVert_{\widetilde{\mathcal{H}}_{k,h}}.$ (9)
Based on this estimator, we can construct the $\kappa$-independent confidence set as follows.
Lemma 3.
For any $\delta\in(0,1)$, define the confidence set as
$\widehat{\mathcal{C}}_{k,h}=\big{\{}\theta\in\Theta\mid\lVert\theta-\widetilde{\theta}_{k,h}\rVert_{\mathcal{H}_{k,h}}\leq\widetilde{\beta}_{k}\big{\}}$
with $\widetilde{\beta}_{k}=\mathcal{O}(\sqrt{d}\log U\log(KH/\delta))$. Then, $\Pr[\theta_{h}^{*}\in\widehat{\mathcal{C}}_{k,h}]\geq 1-\delta,\forall k\in[K],h\in[H]$.
Based on Lemma 3, we can construct the optimistic value function similar to (6) as
$\bar{Q}_{k,h}(s,a)=\bigg{[}r_{h}(s,a)+\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\widetilde{\theta}_{k,h})\bar{V}_{k,h+1}(s^{\prime})+H\widetilde{\beta}_{k}\max_{s^{\prime}\in\mathcal{S}_{h,s,a}}\|\phi_{s,a}^{s^{\prime}}\|_{\mathcal{H}_{k,h}^{-1}}\bigg{]}_{[0,H]}.$ (10)
The detailed algorithm is presented in UCRL-MNL-LL in Algorithm 2. We show the regret guarantee of UCRL-MNL-LL by the following theorem.
Theorem 2.
Suppose Assumption 1 holds. For any $\delta\in(0,1)$, set $\eta=\frac{1}{2}\log(1+U)+(B+1)$, $\lambda=84\sqrt{2}\eta(B+d)$, and $\widetilde{\beta}_{k}=\mathcal{O}(\sqrt{d}\log U\log(KH/\delta))$, with probability at least $1-\delta$, UCRL-MNL-LL algorithm ensures the following regret guarantee
${\operatorname*{Reg}}(K)\leq\widetilde{\mathcal{O}}(\kappa^{-1/2}dH^{2}\sqrt{K}).$
Compared with the result in Theorem 1, Theorem 2 enhances the dependence on $\kappa$ from $\kappa^{-1}$ to $\kappa^{-1/2}$. This is achieved by utilizing local information of the Hessian matrix.
Algorithm 2 UCRL-MNL-LL &\&& UCRL-MNL-LL+
0:  Step size $\eta$, Regularization parameter $\lambda$, confidence width $\widetilde{\beta}_{k}$, confidence parameter $\delta$.
1:  Initialization: $\mathcal{H}_{1,h}=\lambda I,\widehat{\theta}_{1,h}=\mathbf{0}$ for all $h\in[H]$.
2:  for $k=1,\ldots,K$ do
3:     UCRL-MNL-LL: Compute $\bar{Q}_{k,h}(\cdot,\cdot)$ in a backward way as in (10).
4:     UCRL-MNL-LL+: Compute $\widetilde{Q}_{k,h}(\cdot,\cdot)$ in a backward way as in (12).
5:     for $h=1,\ldots,H$ do
6:        UCRL-MNL-LL: Observe $s_{k,h}$, select $a_{k,h}=\operatorname*{arg\,max}_{a\in\mathcal{A}}\bar{Q}_{k,h}(s_{k,h},a)$.
7:        UCRL-MNL-LL+: Observe $s_{k,h}$, select $a_{k,h}=\operatorname*{arg\,max}_{a\in\mathcal{A}}\widetilde{Q}_{k,h}(s_{k,h},a)$.
8:        Update ~$\widetilde{\mathcal{H}}_{k,h}=\mathcal{H}_{k,h}+\eta H_{k,h}(\widetilde{\theta}_{k,h})$.
9:        Compute $\widetilde{\theta}_{k+1,h}=\operatorname*{arg\,min}_{\theta\in\Theta}\langle g_{k,h}(\widetilde{\theta}_{k,h}),\theta-\widetilde{\theta}_{k,h}\rangle+\frac{1}{2\eta}\lVert\theta-\widetilde{\theta}_{k,h}\rVert_{\widetilde{\mathcal{H}}_{k,h}}$ as (8).
10:        Update $\mathcal{H}_{k,h}=\mathcal{H}_{k,h}+H_{k,h}(\widetilde{\theta}_{k+1,h})$.
11:     end for
12:  end for
5.2 Improved Optimistic Value Function Construction
In Section 5.1, we improve the dependence on $\kappa$ by constructing a tighter $\kappa$-independent confidence set. However, the regret of UCRL-MNL-LL in Theorem 2 still have a dependence on $\kappa^{-1/2}$. In this part, we design an improved algorithm that removes the dependence on $\kappa$ in the main term.
By a similar analysis as (7), the regret of UCRL-MNL-LL can be bounded as:
$\operatorname*{Reg}(K)\leq 2H\widetilde{\beta}_{K}\sum_{k=1}^{K}\sum_{h=1}^{H}\max_{s^{\prime}\in\mathcal{S}_{k,h}}\lVert\phi_{k,h}^{s^{\prime}}\rVert_{\mathcal{H}_{k,h}^{-1}}+\sum_{k=1}^{K}\sum_{h=1}^{H}\mathcal{M}_{k,h}.$ (11)
Though $\widetilde{\beta}_{K}$ now is independent of $\kappa$ and $\mathcal{H}_{k,h}^{-1}$ preserves the local information in (11), the norm $\max_{s^{\prime}\in\mathcal{S}_{k,h}}\lVert\phi_{k,h}^{s^{\prime}}\rVert_{\mathcal{H}_{k,h}^{-1}}$ is still in a global manner due to the maximum operation, leading to a $\kappa^{-1/2}$ dependence. To address this issue, we propose a new construction of the optimistic value function through a more refined analysis. Specifically, we employ a second-order Taylor expansion to more accurately bound the value difference arising from the estimation error in the following lemma.
Lemma 4.
Suppose Lemma 3 holds. For any function $V:\mathcal{S}\to[0,H]$ and $(h,s,a)\in[H]\times\mathcal{S}\times\mathcal{A}$, with probability at least $1-\delta$, it holds that
$\bigg{|}{\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\widetilde{\theta}_{k,h})V(s^{\prime})-\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\theta_{h}^{*})V(s^{\prime})}\bigg{|}\leq\epsilon_{s,a}^{\mathtt{fst}}+\epsilon_{s,a}^{\mathtt{snd}},$
where
$\epsilon_{s,a}^{\mathtt{fst}}=H\widetilde{\beta}_{k}\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\widetilde{\theta}_{k,h})\Big{\lVert}\phi_{s,a}^{s^{\prime}}-\sum_{s^{\prime\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime\prime}}(\widetilde{\theta}_{k,h})\phi_{s,a}^{s^{\prime\prime}}\Big{\rVert}_{\mathcal{H}_{k,h}^{-1}},\quad\epsilon_{s,a}^{\mathtt{snd}}=\frac{5}{2}H\widetilde{\beta}_{k}^{2}\max_{s^{\prime}\in\mathcal{S}_{h,s,a}}\lVert\phi_{s,a}^{s^{\prime}}\rVert_{\mathcal{H}_{k,h}^{-1}}^{2}.$
Based on Lemma 4, we construct the optimistic value function as follows:
$\widetilde{Q}_{k,h}(s,a)=\bigg{[}r_{h}(s,a)+\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\widetilde{\theta}_{k,h})\widetilde{V}_{k,h+1}(s^{\prime})+\epsilon_{s,a}^{\mathtt{fst}}+\epsilon_{s,a}^{\mathtt{snd}}\bigg{]}_{[0,H]}.$ (12)
The detailed algorithm is presented in UCRL-MNL-LL+ in Algorithm 2. We show the regret guarantee of UCRL-MNL-LL+ in the following theorem.
Theorem 3.
Suppose Assumption 1 holds. For any $\delta\in(0,1)$, set $\eta=\frac{1}{2}\log(1+U)+(B+1)$, $\lambda=84\sqrt{2}\eta(B+d)$ and $\widetilde{\beta}_{k}=\mathcal{O}(\sqrt{d}\log U\log(KH/\delta))$, with probability at least $1-\delta$, UCRL-MNL-LL+ algorithm ensures the following regret guarantee
${\operatorname*{Reg}}(K)\leq\widetilde{\mathcal{O}}\big{(}dH^{2}\sqrt{K}+\kappa^{-1}d^{2}H^{2}\big{)}.$
Remark 3.
Theorem 3 moves the exponentially small $\kappa$-dependence to the lower-order term, which is independent of $K$. Compared with the optimal regret of $\widetilde{\mathcal{O}}(d\sqrt{H^{3}K})$ for linear mixture MDPs (Zhou et al., 2021) and linear MDPs (He et al., 2023), the higher-order term only differs by a factor of $H^{1/2}$, almost matching the same computational and statistical efficiency simultaneously.
6 Lower Bound
In this section, we establish the first lower bound for MNL mixture MDPs by presenting a novel reduction, which connects MNL mixture MDPs and logistic bandits.
Consider the following logistic bandit problem (Faury et al., 2020): at each round $k\in[K]$, the learner selects an action $x_{k}\in\mathcal{X}$ and receives a reward $r_{k}$ sampled from Bernoulli distribution with mean $\mu(x^{\top}\theta^{*})=(1+\exp(-x^{\top}\theta^{*}))^{-1}$, where $\theta^{*}\in\{\theta\in\mathbb{R}^{d}\mid\lVert\theta\rVert_{2}\leq B\}$ is the unknown parameter. The learner aims to minimize the regret: ${\operatorname*{Reg}}^{\mathtt{LB}}(K)=\max_{x\in\mathcal{X}}\sum_{k=1}^{K}\mu(x^{\top}\theta^{*})-\sum_{k=1}^{K}\mu(x_{k}^{\top}\theta^{*})$.
Theorem 4.
For any logistic bandit $\mathcal{B}$, there exists an MNL mixture MDP $\mathcal{M}$ such that learning $\mathcal{M}$ is as hard as learning $H/2$ independent instances of $\mathcal{B}$ simultaneously.
Corollary 1.
For any problem instance $\{\theta_{h}^{*}\}_{h=1}^{H}$ and for $K\geq d^{2}\kappa^{*}$, there exists an MNL mixture MDP with infinite action space such that $\operatorname*{Reg}(K)\geq\Omega(dH\sqrt{K\kappa^{*}})$.
Remark 4.
Corollary 1 can be proved by combining Theorem 4 and the $\Omega(d\sqrt{K\kappa^{*}})$ lower bound for logistic bandits with infinite arms by Abeille et al. (2021). To our knowledge, a lower bound for logistic bandits with finite arms has not been established, which is beyond the scope of this work. Consequently, it remains open for the lower bound of MNL mixture MDPs with finite action space.
Remark 5.
While Corollary 1 is not a strict lower bound for MNL mixture MDPs with finite action space, it gives a basis to conjecture our results may be optimal in $d$ and $K$.
Dependence on $H$.
MNL mixture MDPs can be viewed as a generalization of linear mixture MDPs (Ayoub et al., 2020; Zhou et al., 2021). The pioneering work by Ayoub et al. (2020) achieved an $\widetilde{\mathcal{O}}(dH^{2}\sqrt{K})$ regret for linear mixture MDPs, which matches our results in Theorem 3, differing only on the lower-order term. Zhou et al. (2021) enhanced the dependence on $H$ and attained an optimal regret of $\widetilde{\mathcal{O}}(d\sqrt{H^{3}K})$. This was made possible by recognizing that the value function in linear mixture MDPs is also linear, allowing for direct learning of value functions while incorporating their variance information. In contrast, the value function does not conform to a logistic model for MNL mixture MDPs, posing a significant challenge in using variance information of value functions. It remains open whether similar improvements in $H$ are attainable for MNL mixture MDPs.
Dependence on $\kappa$.
Though we fully remove the $\kappa$ dependence in Theorem 3 in the higher-order term and move it to the lower order term, Corollary 1 suggests that there may still be some room for improvement in the dependency on $\kappa$. The $\widetilde{\mathcal{O}}(d\sqrt{K\kappa^{*}})$ dependence has been established in the binary logistic bandit problem by Abeille et al. (2021). However, their techniques are limited to the binary case and are not directly applicable to MNL mixture MDPs which have multinomial rewards. For the multinomial feedback settings, Périvier and Goyal (2022) studied the MNL contextual bandit problem and achieved a regret bound of $\widetilde{\mathcal{O}}(d\sqrt{K\kappa^{*}}+\kappa^{-1})$. However, their results only hold for the scenarios with uniform revenue, where each choice yields identical rewards. In contrast, MNL mixture MDPs are more challenging as they involve diverse revenue values, represented by the value functions of different states. Thus, their techniques also can not be applied to MNL mixture MDPs. It remains open whether the $\widetilde{\mathcal{O}}(dH\sqrt{K\kappa^{*}})$ result can be achieved for MNL mixture MDPs.
7 Conclusion and Future Work
In this work, we address both the computational and statistical challenges for MDPs with multinomial logit function approximation. We first propose an algorithm that attains the previously best-known $\widetilde{\mathcal{O}}(\kappa^{-1}dH^{2}\sqrt{K})$ regret with $\mathcal{O}(1)$ computational and storage costs per episode. Next, we identify that the statistical efficiency can be further improved by exploiting the local information. We propose two algorithms that not only maintain $\mathcal{O}(1)$ computational and storage costs but achieve improved regrets of $\widetilde{\mathcal{O}}(\kappa^{-1/2}dH^{2}\sqrt{K})$ and $\widetilde{\mathcal{O}}(dH^{2}\sqrt{K}+\kappa^{-1}d^{2}H^{2})$, respectively. We also establish the first lower bound, justifying the optimality of our results in $d$ and $K$. To the best of our knowledge, this is the first work that achieves almost the same computational and statistical efficiency as linear function approximation while employing non-linear function approximation.
There are several interesting directions for future work. First, there still exists a gap between the upper and lower bounds and how to close this gap is an open problem. Besides, while this work focuses on stochastic rewards, extending this model to the non-stationary settings and studying the dynamic regret (Wei and Luo, 2021; Fei et al., 2020; Li et al., 2023) is another interesting direction.
Appendix A Properties of the Multinomial Logit Function
This section collects several key properties of the multinomial logit function.
Without loss of generality, we assume $\forall\mathcal{S}_{h,s,a},\exists s^{\prime}\in\mathcal{S}_{h,s,a}$ such that $\phi(s^{\prime}\>|\>s,a)=\mathbf{0}$. Otherwise, we can always define a new feature mapping $\phi^{\prime}(s^{\prime\prime}\>|\>s,a)=\phi(s^{\prime}\>|\>s,a)-\phi(s^{\prime\prime}\>|\>s,a)$ for any $s^{\prime\prime}\in\mathcal{S}_{h,s,a}$ such that $\phi^{\prime}(s^{\prime}\>|\>s,a)=\mathbf{0}$ and the transition induced by $\phi^{\prime}$ is the same as that induced by $\phi$. We denote this state as $\dot{s}_{h,s,a}$ and $\dot{\mathcal{S}}_{h,s,a}=\mathcal{S}_{h,s,a}\backslash\{\dot{s}_{h,s,a}\}$.
Recall the following definitions in the paper:
$f_{k,h}(\theta)$ $=-\sum_{s^{\prime}\in\mathcal{S}_{k,h}}y_{k,h}^{s^{\prime}}\log p_{k,h}^{s^{\prime}}(\theta),\quad g_{k,h}(\theta)=\sum_{s^{\prime}\in\mathcal{S}_{k,h}}(p_{k,h}^{s^{\prime}}(\theta)-y_{k,h}^{s^{\prime}})\phi_{k,h}^{s^{\prime}}$
$H_{k,h}(\theta)$ $=\sum_{s^{\prime}\in\mathcal{S}_{k,h}}p_{k,h}^{s^{\prime}}(\theta)\phi_{k,h}^{s^{\prime}}(\phi_{k,h}^{s^{\prime}})^{\top}-\sum_{s^{\prime}\in\mathcal{S}_{k,h}}\sum_{s^{\prime\prime}\in\mathcal{S}_{k,h}}p_{k,h}^{s^{\prime}}(\theta)p_{k,h}^{s^{\prime\prime}}(\theta)\phi_{k,h}^{s^{\prime}}(\phi_{k,h}^{s^{\prime\prime}})^{\top}.$
Lemma 5.
The following statements hold for any $k\in[K],h\in[H]$:
$H_{k,h}(\theta)\succeq\sum_{s^{\prime}\in\dot{\mathcal{S}}_{k,h}}p_{k,h}^{s^{\prime}}(\theta)p_{k,h}^{\dot{s}_{k,h}}(\theta)\phi_{k,h}^{s^{\prime}}(\phi_{k,h}^{s^{\prime}})^{\top}\succeq\kappa\sum_{s^{\prime}\in\dot{\mathcal{S}}_{k,h}}\phi_{k,h}^{s^{\prime}}(\phi_{k,h}^{s^{\prime}})^{\top}.$
Proof.
Note that
$\forall x,y\in\mathbb{R}^{d},(x-y)(x-y)^{\top}=xx^{\top}+yy^{\top}-xy^{\top}-yx^{\top}\succeq 0\Longrightarrow xx^{\top}+yy^{\top}\succeq xy^{\top}+yx^{\top}.$
Then, we have
$H_{k,h}(\theta)$
=\ = $\sum_{s^{\prime}\in\mathcal{S}_{k,h}}p_{k,h}^{s^{\prime}}(\theta)\phi_{k,h}^{s^{\prime}}(\phi_{k,h}^{s^{\prime}})^{\top}-\sum_{s^{\prime}\in\mathcal{S}_{k,h}}\sum_{s^{\prime\prime}\in\mathcal{S}_{k,h}}p_{k,h}^{s^{\prime}}(\theta)p_{k,h}^{s^{\prime\prime}}(\theta)\phi_{k,h}^{s^{\prime}}(\phi_{k,h}^{s^{\prime\prime}})^{\top}$
=\ = $\sum_{s^{\prime}\in\dot{\mathcal{S}}_{k,h}}p_{k,h}^{s^{\prime}}(\theta)\phi_{k,h}^{s^{\prime}}(\phi_{k,h}^{s^{\prime}})^{\top}-\sum_{s^{\prime}\in\dot{\mathcal{S}}_{k,h}}\sum_{s^{\prime\prime}\in\dot{\mathcal{S}}_{k,h}}p_{k,h}^{s^{\prime}}(\theta)p_{k,h}^{s^{\prime\prime}}(\theta)\phi_{k,h}^{s^{\prime}}(\phi_{k,h}^{s^{\prime\prime}})^{\top}$
=\ = $\sum_{s^{\prime}\in\dot{\mathcal{S}}_{k,h}}p_{k,h}^{s^{\prime}}(\theta)\phi_{k,h}^{s^{\prime}}(\phi_{k,h}^{s^{\prime}})^{\top}-\frac{1}{2}\sum_{s^{\prime}\in\dot{\mathcal{S}}_{k,h}}\sum_{s^{\prime\prime}\in\dot{\mathcal{S}}_{k,h}}p_{k,h}^{s^{\prime}}(\theta)p_{k,h}^{s^{\prime\prime}}(\theta)\left(\phi_{k,h}^{s^{\prime}}(\phi_{k,h}^{s^{\prime\prime}})^{\top}+\phi_{k,h}^{s^{\prime\prime}}(\phi_{k,h}^{s^{\prime}})^{\top}\right)$
$\succeq$ $\sum_{s^{\prime}\in\dot{\mathcal{S}}_{k,h}}p_{k,h}^{s^{\prime}}(\theta)\phi_{k,h}^{s^{\prime}}(\phi_{k,h}^{s^{\prime}})^{\top}-\frac{1}{2}\sum_{s^{\prime}\in\dot{\mathcal{S}}_{k,h}}\sum_{s^{\prime\prime}\in\dot{\mathcal{S}}_{k,h}}p_{k,h}^{s^{\prime}}(\theta)p_{k,h}^{s^{\prime\prime}}(\theta)\left(\phi_{k,h}^{s^{\prime}}(\phi_{k,h}^{s^{\prime}})^{\top}+\phi_{k,h}^{s^{\prime\prime}}(\phi_{k,h}^{s^{\prime\prime}})^{\top}\right)$
=\ = $\sum_{s^{\prime}\in\dot{\mathcal{S}}_{k,h}}p_{k,h}^{s^{\prime}}(\theta)\phi_{k,h}^{s^{\prime}}(\phi_{k,h}^{s^{\prime}})^{\top}-\sum_{s^{\prime}\in\dot{\mathcal{S}}_{k,h}}\sum_{s^{\prime\prime}\in\dot{\mathcal{S}}_{k,h}}p_{k,h}^{s^{\prime}}(\theta)p_{k,h}^{s^{\prime\prime}}(\theta)\phi_{k,h}^{s^{\prime}}(\phi_{k,h}^{s^{\prime}})^{\top}$
=\ = $\sum_{s^{\prime}\in\dot{\mathcal{S}}_{k,h}}p_{k,h}^{s^{\prime}}(\theta)\Big{(}1-\sum_{s^{\prime\prime}\in\dot{\mathcal{S}}_{k,h}}p_{k,h}^{s^{\prime\prime}}(\theta)\Big{)}\phi_{k,h}^{s^{\prime}}(\phi_{k,h}^{s^{\prime}})^{\top}$
=\ = $\sum_{s^{\prime}\in\dot{\mathcal{S}}_{k,h}}p_{k,h}^{s^{\prime}}(\theta)p_{k,h}^{\dot{s}_{k,h}}(\theta)\phi_{k,h}^{s^{\prime}}(\phi_{k,h}^{s^{\prime}})^{\top}$
[…]
where the last inequality holds by the definition of $\kappa$ in Assumption 1. This finishes the proof. ∎
Lemma 6.
Suppose $\lambda\geq 1$, define $\widetilde{\phi}_{k,h}^{s^{\prime}}=\phi_{k,h}^{s^{\prime}}-\sum_{s^{\prime\prime}\in\mathcal{S}_{k,h}}p_{k,h}^{s^{\prime\prime}}(\widetilde{\theta}_{k+1,h})\phi_{k,h}^{s^{\prime\prime}}$, for any $k\in[K],h\in[H]$, the following statements hold:
$(1)\quad\sum_{i=1}^{k}\max_{s^{\prime}\in\mathcal{S}_{i,h}}\lVert\phi_{i,h}^{s^{\prime}}\rVert_{\widehat{\Sigma}_{i,h}^{-1}}^{2}\leq\frac{4}{\kappa}d\log\left(1+\frac{kU}{\lambda d}\right)$
$(2)\quad\sum_{i=1}^{k}\sum_{s^{\prime}\in\mathcal{S}_{i,h}}p_{i,h}^{s^{\prime}}(\widetilde{\theta}_{i+1,h})p_{i,h}^{\dot{s}_{k,h}}(\widetilde{\theta}_{i+1,h})\lVert\phi_{i,h}^{s^{\prime}}\rVert_{\mathcal{H}_{i,h}^{-1}}^{2}\leq 2d\log\left(1+\frac{k}{\lambda d}\right)$
$(3)\quad\sum_{i=1}^{k}\max_{s^{\prime}\in\mathcal{S}_{i,h}}\lVert\phi_{i,h}^{s^{\prime}}\rVert_{\mathcal{H}_{i,h}^{-1}}^{2}\leq\frac{2}{\kappa}d\log\left(1+\frac{k}{\lambda d}\right)$
$(4)\quad\sum_{i=1}^{k}\sum_{s^{\prime}\in\mathcal{S}_{i,h}}p_{i,h}^{s^{\prime}}(\widetilde{\theta}_{i+1,h})\lVert\widetilde{\phi}_{i,h}^{s^{\prime}}\rVert_{\mathcal{H}_{i,h}^{-1}}^{2}\leq 2d\log\left(1+\frac{k}{\lambda d}\right)$
$(5)\quad\sum_{i=1}^{k}\max_{s^{\prime}\in\mathcal{S}_{i,h}}\lVert\widetilde{\phi}_{i,h}^{s^{\prime}}\rVert_{\mathcal{H}_{i,h}^{-1}}^{2}\leq\frac{2}{\kappa}d\log\left(1+\frac{k}{\lambda d}\right)$
Proof.
We prove the three statements individually.
Proof of statement (1).
By the definition that $\widehat{\Sigma}_{k+1,h}=\widehat{\Sigma}_{k,h}+\frac{\kappa}{2}\sum_{s^{\prime}\in\mathcal{S}_{k,h}}\phi_{k,h}^{s^{\prime}}(\phi_{k,h}^{s^{\prime}})^{\top}$, we have
$\det(\widehat{\Sigma}_{k+1,h})=\det(\widehat{\Sigma}_{k,h})\Big{(}1+\frac{\kappa}{2}\sum_{s^{\prime}\in\mathcal{S}_{k,h}}\lVert\phi_{k,h}^{s^{\prime}}\rVert^{2}_{\widehat{\Sigma}_{k,h}^{-1}}\Big{)}.$
Then, we get
$\sum_{i=1}^{k}\log\Big{(}1+\frac{\kappa}{2}\sum_{s^{\prime}\in\mathcal{S}_{i,h}}\lVert\phi_{i,h}^{s^{\prime}}\rVert^{2}_{\widehat{\Sigma}_{i,h}^{-1}}\Big{)}\leq\log\left(\frac{\det(\widehat{\Sigma}_{k+1,h})}{\det(\widehat{\Sigma}_{1,h})}\right)\leq d\log\left(1+\frac{kU}{\lambda d}\right),$ (13)
where the last inequality holds by determinant trace inequality in Lemma 16. Since $\lambda\geq 1$, we have $\max_{s^{\prime}\in\mathcal{S}_{i,h}}\lVert\phi_{i,h}^{s^{\prime}}\rVert_{\widehat{\Sigma}_{i,h}^{-1}}\leq 1$, thus, we have
$\sum_{i=1}^{k}\max_{s^{\prime}\in\mathcal{S}_{i,h}}\lVert\phi_{i,h}^{s^{\prime}}\rVert_{\widehat{\Sigma}_{i,h}^{-1}}$ $\leq\frac{2}{\kappa}\sum_{i=1}^{k}\min\bigg{\{}1,\frac{\kappa}{2}\sum_{s^{\prime}\in\mathcal{S}_{i,h}}\lVert\phi_{i,h}^{s^{\prime}}\rVert_{\widehat{\Sigma}_{i,h}^{-1}}\bigg{\}}$
$\leq\frac{4}{\kappa}\sum_{i=1}^{k}\log\Bigg{(}1+\min\bigg{\{}1,\frac{\kappa}{2}\sum_{s^{\prime}\in\mathcal{S}_{i,h}}\lVert\phi_{i,h}^{s^{\prime}}\rVert_{\widehat{\Sigma}_{i,h}^{-1}}\bigg{\}}\Bigg{)}$
$\leq\frac{4}{\kappa}\sum_{i=1}^{k}\log\Big{(}1+\frac{\kappa}{2}\sum_{s^{\prime}\in\mathcal{S}_{i,h}}\lVert\phi_{i,h}^{s^{\prime}}\rVert^{2}_{\widehat{\Sigma}_{i,h}^{-1}}\Big{)}$
$\leq\frac{4}{\kappa}d\log\left(1+\frac{kU}{\lambda d}\right),$
where the first inequality holds by the fact that $z\leq 2\log(1+z)$ for any $z\in[0,1]$ and the last inequality holds by Equation (13).
Proof of statement (2).
By Lemma 5, we have $H_{k,h}(\theta)\succeq\sum_{s^{\prime}\in\dot{\mathcal{S}}_{k,h}}p_{k,h}^{s^{\prime}}(\theta)p_{k,h}^{\dot{s}_{k,h}}(\theta)\phi_{k,h}^{s^{\prime}}(\phi_{k,h}^{s^{\prime}})^{\top}$. Thus, we have
$\mathcal{H}_{k+1,h}\succeq\mathcal{H}_{k,h}+\sum_{s^{\prime}\in\dot{\mathcal{S}}_{k,h}}p_{k,h}^{s^{\prime}}(\widetilde{\theta}_{k+1,h})p_{k,h}^{\dot{s}_{k,h}}(\widetilde{\theta}_{k+1,h})\phi_{k,h}^{s^{\prime}}(\phi_{k,h}^{s^{\prime}})^{\top}$
Then, we get
$\det(\mathcal{H}_{i+1,h})\geq\det(\mathcal{H}_{i,h})\Big{(}1+\sum_{s^{\prime}\in\dot{\mathcal{S}}_{i,h}}p_{i,h}^{s^{\prime}}(\widetilde{\theta}_{i+1,h})p_{i,h}^{\dot{s}_{i,h}}(\widetilde{\theta}_{i+1,h})\lVert\phi_{i,h}^{s^{\prime}}\rVert^{2}_{\mathcal{H}_{i,h}^{-1}}\Big{)}.$
Since $\lambda\geq 1$, we have $\sum_{s^{\prime}\in\dot{\mathcal{S}}_{i,h}}p_{i,h}^{s^{\prime}}(\widetilde{\theta}_{i+1,h})p_{i,h}^{\dot{s}_{i,h}}(\widetilde{\theta}_{i+1,h})\lVert\phi_{i,h}^{s^{\prime}}\rVert^{2}_{\mathcal{H}_{i,h}^{-1}}\leq 1$. Using the fact that $z\leq 2\log(1+z)$ for any $z\in[0,1]$, we get
$\sum_{i=1}^{k}\sum_{s^{\prime}\in\dot{\mathcal{S}}_{i,h}}p_{i,h}^{s^{\prime}}(\widetilde{\theta}_{i+1,h})p_{i,h}^{\dot{s}_{i,h}}(\widetilde{\theta}_{i+1,h})\lVert\phi_{i,h}^{s^{\prime}}\rVert^{2}_{\mathcal{H}_{i,h}^{-1}}$
[…]
Proof of statement (3).
By Lemma 5, we have $\mathcal{H}_{k+1,h}\succeq\mathcal{H}_{k+1,h}+\kappa\sum_{s^{\prime}\in\dot{\mathcal{S}}_{k,h}}\phi_{k,h}^{s^{\prime}}(\phi_{k,h}^{s^{\prime}})^{\top}$. Since $\lambda\geq 1$, we have $\kappa\max_{s^{\prime}\in\mathcal{S}_{i,h}}\lVert\phi_{i,h}^{s^{\prime}}\rVert_{\mathcal{H}_{i,h}^{-1}}\leq\kappa$. Using the fact that $z\leq 2\log(1+z)$ for any $z\in[0,1]$. By a similar analysis as the statement (2), we have
$\sum_{i=1}^{k}\max_{s^{\prime}\in\mathcal{S}_{i,h}}\lVert\phi_{i,h}^{s^{\prime}}\rVert_{\mathcal{H}_{i,h}^{-1}}^{2}$ $\leq\frac{2}{\kappa}\sum_{i=1}^{k}\log\Big{(}1+\max_{s^{\prime}\in\mathcal{S}_{i,h}}\lVert\phi_{i,h}^{s^{\prime}}\rVert_{\mathcal{H}_{i,h}^{-1}}\Big{)}$
$\leq\frac{2}{\kappa}\log\left(\frac{\det(\mathcal{H}_{k+1,h})}{\det(\mathcal{H}_{1,h})}\right)$
$\leq\frac{2}{\kappa}d\log\left(1+\frac{k}{\lambda d}\right).$
Proof of statement (4).
By the definition of $H_{k,h}(\theta)$, we have
$H_{i,h}(\theta)$ $=\sum_{s^{\prime}\in\mathcal{S}_{i,h}}p_{i,h}^{s^{\prime}}(\theta)\phi_{i,h}^{s^{\prime}}(\phi_{i,h}^{s^{\prime}})^{\top}-\sum_{s^{\prime}\in\mathcal{S}_{i,h}}\sum_{s^{\prime\prime}\in\mathcal{S}_{i,h}}p_{i,h}^{s^{\prime}}(\theta)p_{i,h}^{s^{\prime\prime}}(\theta)\phi_{i,h}^{s^{\prime}}(\phi_{i,h}^{s^{\prime\prime}})^{\top}$
$=\mathbb{E}_{s^{\prime}\in p_{i,h}(\theta)}[\phi_{i,h}^{s^{\prime}}(\phi_{i,h}^{s^{\prime}})^{\top}]-\mathbb{E}_{s^{\prime}\in p_{i,h}(\theta)}[\phi_{i,h}^{s^{\prime}}]\big{(}\mathbb{E}_{s^{\prime}\in p_{i,h}(\theta)}[\phi_{i,h}^{s^{\prime}}]\big{)}^{\top}$
$=\mathbb{E}_{s^{\prime}\in p_{i,h}(\theta)}\big{[}(\phi_{i,h}^{s^{\prime}}-\mathbb{E}_{s^{\prime\prime}\in p_{i,h}(\theta)}\phi_{i,h}^{s^{\prime\prime}})(\phi_{i,h}^{s^{\prime}}-\mathbb{E}_{s^{\prime\prime}\in p_{i,h}(\theta)}\phi_{i,h}^{s^{\prime\prime}})^{\top}\big{]}$
Thus, we have
$H_{i,h}(\widetilde{\theta}_{i+1,h})\succeq\sum_{s^{\prime}\in\mathcal{S}_{i,h}}p_{i,h}(\widetilde{\theta}_{i+1,h})(\widetilde{\phi}_{i,h}^{s^{\prime}})(\widetilde{\phi}_{i,h}^{s^{\prime}})^{\top}.$ (14)
Then, we get
$\mathcal{H}_{k+1,h}\succeq\mathcal{H}_{k,h}+\sum_{s^{\prime}\in\mathcal{S}_{k,h}}p_{k,h}(\widetilde{\theta}_{k+1,h})(\widetilde{\phi}_{k,h}^{s^{\prime}})(\widetilde{\phi}_{k,h}^{s^{\prime}})^{\top}$
The remaining proof is similar to the proof of statement (2).
Proof of statement (5).
By (14), we have
$H_{i,h}(\widetilde{\theta}_{i+1,h})\succeq\sum_{s^{\prime}\in\mathcal{S}_{i,h}}p_{i,h}(\widetilde{\theta}_{i+1,h})(\widetilde{\phi}_{i,h}^{s^{\prime}})(\widetilde{\phi}_{i,h}^{s^{\prime}})^{\top}\succeq\kappa\sum_{s^{\prime}\in\mathcal{S}_{i,h}}(\widetilde{\phi}_{i,h}^{s^{\prime}})(\widetilde{\phi}_{i,h}^{s^{\prime}})^{\top}.$
Then, the remaining proof is similar to the proof of statement (3). ∎
Appendix B Proof of Lemma 1
B.1 Main Proof
Proof.
We define
$\bar{f}_{k,h}(\theta)=\mathbb{E}_{y_{k,h}}[f_{k,h}(\theta)\mid\mathcal{F}_{k,h}],\quad\bar{g}_{k,h}(\theta)=\mathbb{E}_{y_{k,h}}[g_{k,h}(\theta)\mid\mathcal{F}_{k,h}].$
By Property 1, we have $f_{k,h}(\theta)$ is exponential concave such that
$f_{k,h}(\widehat{\theta}_{k,h})\leq f_{k,h}(\theta_{h}^{*})+\langle g_{k,h}(\theta_{h}^{*}),\widehat{\theta}_{k,h}-\theta_{h}^{*}\rangle-\frac{\kappa}{2}\lVert(\widehat{\theta}_{k,h}-\theta_{h}^{*})\rVert_{W_{k,h}}.$
Taking expectations on both sides, we have
$\bar{f}_{k,h}(\widehat{\theta}_{k,h})\leq\bar{f}_{k,h}(\theta_{h}^{*})+\langle\bar{g}_{k,h}(\theta_{h}^{*}),\widehat{\theta}_{k,h}-\theta_{h}^{*}\rangle-\frac{\kappa}{2}\lVert(\widehat{\theta}_{k,h}-\theta_{h}^{*})\rVert_{W_{k,h}}.$ (15)
Based the property of KL diverge, we ensure the true parameter $\theta_{h}^{*}$ is the minimizer of the expected loss function by the following lemma.
Lemma 7.
For any $\theta\in\Theta$, we have $\bar{f}_{k,h}(\theta_{h}^{*})\leq\bar{f}_{k,h}(\theta)$.
Combining Lemma 7 and Equation (15), we have
00 0 $\leq\bar{f}_{k,h}(\widehat{\theta}_{k,h})-\bar{f}_{k,h}(\theta_{h}^{*})$
$\leq\langle\bar{g}_{k,h}(\widehat{\theta}_{k,h}),\widehat{\theta}_{k,h}-\theta_{h}^{*}\rangle-\frac{\kappa}{2}\lVert\theta_{h}^{*}-\widehat{\theta}_{k,h}\rVert_{W_{k,h}}^{2}$
$\leq\langle g_{k,h}(\widehat{\theta}_{k,h}),\widehat{\theta}_{k,h}-\theta_{h}^{*}\rangle-\frac{\kappa}{2}\lVert\theta_{h}^{*}-\widehat{\theta}_{k,h}\rVert_{W_{k,h}}^{2}+\langle\bar{g}_{k,h}(\widehat{\theta}_{k,h})-g_{k,h}(\widehat{\theta}_{k,h}),\widehat{\theta}_{k,h}-\theta_{h}^{*}\rangle.$ (16)
By standard analysis of OMD in Lemma 17, it holds
$2g_{k,h}(\widehat{\theta}_{k,h})(\widehat{\theta}_{k,h}-\theta_{h}^{*})\leq\lVert g_{k,h}(\widehat{\theta}_{k,h})\rVert_{\widehat{\Sigma}_{k+1,h}}+\lVert\widehat{\theta}_{k,h}-\theta_{h}^{*}\rVert_{\widehat{\Sigma}_{k+1,h}}-\lVert\widehat{\theta}_{k+1,h}-\theta_{h}^{*}\rVert_{\widehat{\Sigma}_{k+1,h}}.$ (17)
Combining Equation (16) and Equation (17), we have
00 0 $\leq\frac{1}{2}\lVert g_{k,h}(\widehat{\theta}_{k,h})\rVert_{\widehat{\Sigma}_{k+1,h}^{-1}}^{2}+\frac{1}{2}\lVert\widehat{\theta}_{k,h}-\theta_{h}^{*}\rVert_{\widehat{\Sigma}_{k+1,h}}^{2}-\frac{1}{2}\lVert\widehat{\theta}_{k+1,h}-\theta_{h}^{*}\rVert_{\widehat{\Sigma}_{k+1,h}}^{2}$
$\hskip 85.35826pt-\frac{\kappa}{2}\lVert\theta_{h}^{*}-\widehat{\theta}_{k,h}\rVert_{W_{k,h}}^{2}+\langle\bar{g}_{k,h}(\widehat{\theta}_{k,h})-g_{k,h}(\widehat{\theta}_{k,h}),\widehat{\theta}_{k,h}-\theta_{h}^{*}\rangle.$ (18)
First, we consider the first term . We show the gradient can be bounded by the following lemma.
Lemma 8.
For any positive semi-definite matrix $Z$, it holds that
$\lVert g_{k,h}(\theta)\rVert_{Z}^{2}\leq 4\max_{s^{\prime}\in\mathcal{S}_{k,h}}\lVert\phi_{k,h}^{s^{\prime}}\rVert_{Z}^{2}.$ (19)
Then, we consider the second term. Note that $\widehat{\Sigma}_{k+1,h}=\widehat{\Sigma}_{k,h}+\frac{\kappa}{2}W_{k,h}$, we have
$\lVert\widehat{\theta}_{k,h}-\theta^{*}\rVert_{\widehat{\Sigma}_{k+1,h}}^{2}=\lVert\widehat{\theta}_{k,h}-\theta^{*}\rVert_{\widehat{\Sigma}_{k,h}}^{2}+\frac{\kappa}{2}\lVert\widehat{\theta}_{k,h}-\theta^{*}\rVert_{W_{k,h}}^{2}.$ (20)
Next, we bound $\langle\bar{g}_{k,h}(\widehat{\theta}_{k,h})-g_{k,h}(\widehat{\theta}_{k,h}),\widehat{\theta}_{k,h}-\theta_{h}^{*}\rangle$, which is a martingale difference sequence.
Lemma 9.
For any $\delta\in(0,1)$ and $\theta_{1},\ldots,\theta_{k},\theta^{*}\in[0,B]^{d}$, with probability at least $1-\delta$, for any $k\in[K],h\in[H]$ it holds that
$\sum_{i=1}^{k}\langle\bar{g}_{i,h}(\theta_{i})-g_{i,h}(\theta_{i}),\theta_{i}-\theta_{h}^{*}\rangle\leq\frac{\kappa}{4}\sum_{i=1}^{k}\lVert\theta_{i}-\theta^{*}\rVert_{W_{i,h}}^{2}+(\frac{4}{\kappa}+8B)\gamma_{k}.$ (21)
where $\gamma_{k}=\log\frac{2k^{2}H\log(kU)}{\delta}$.
Combining (18), (19), and (20), we have
$\lVert\widehat{\theta}_{k+1,h}-\theta_{h}^{*}\rVert_{\widehat{\Sigma}_{k+1,h}}^{2}$ $\leq\lVert\widehat{\theta}_{k,h}-\theta_{h}^{*}\rVert_{\widehat{\Sigma}_{k+1,h}}^{2}-\kappa\lVert\theta_{h}^{*}-\widehat{\theta}_{k,h}\rVert_{W_{k,h}}^{2}+\lVert g_{k,h}(\widehat{\theta}_{k,h})\rVert_{\widehat{\Sigma}_{k+1,h}^{-1}}^{2}$
$\hskip 28.45274pt+2\langle\bar{g}_{k,h}(\widehat{\theta}_{k,h})-g_{k,h}(\widehat{\theta}_{k,h}),\widehat{\theta}_{k,h}-\theta_{h}^{*}\rangle$
$\leq\lVert\widehat{\theta}_{k,h}-\theta_{h}^{*}\rVert_{\widehat{\Sigma}_{k,h}}^{2}+4\max_{s^{\prime}\in\mathcal{S}_{k,h}}\lVert\phi_{k,h}^{s^{\prime}}\rVert_{\widehat{\Sigma}_{k+1,h}^{-1}}^{2}+(\frac{8}{\kappa}+16B)\gamma_{k}$ (22)
where the first inequality holds by rearranging the terms in (18), the second holds by (19), (20) and (21). Summing (22) from $i=1$ to $k$, we have
$\lVert\widehat{\theta}_{k+1,h}-\theta_{h}^{*}\rVert_{\widehat{\Sigma}_{k+1,h}}^{2}$ $\leq\lambda B^{2}+4\sum_{i=1}^{k}\max_{s^{\prime}\in\mathcal{S}_{i,h}}\lVert\phi_{i,h}^{s^{\prime}}\rVert_{\widehat{\Sigma}_{i+1,h}^{-1}}^{2}+(\frac{8}{\kappa}+16B)\gamma_{k}$
$\leq\lambda B^{2}+\frac{8}{\kappa}d\log\left(1+\frac{kU}{\lambda d}\right)+(\frac{8}{\kappa}+16B)\gamma_{k},$
where the second inequality holds by $\lVert\phi_{i,h}^{s^{\prime}}\rVert_{2}\leq 1$ and $\widehat{\Sigma}_{i,h}\geq I,\forall i\in[K]$, the last inequality is by the elliptical potential in Lemma 6. This finishes the proof. ∎
B.2 Proof of Auxiliary Lemmas
In this section, we provide the proofs of the lemmas used in Appendix B.1.
B.2.1 Proof of Lemma 7
Proof.
By the definition of $\bar{f}_{k,h}(\theta)$, we have
$\bar{f}_{k,h}(\theta)-\bar{f}_{k,h}(\theta_{h}^{*})$ $=-\sum_{s^{\prime}\in\mathcal{S}_{k,h}}p_{k,h}^{s^{\prime}}(\theta_{h}^{*})\log p_{k,h}^{s^{\prime}}(\theta)+\sum_{s^{\prime}\in\mathcal{S}_{k,h}}p_{k,h}^{s^{\prime}}(\theta_{h}^{*})\log p_{k,h}^{s^{\prime}}(\theta_{h}^{*})\geq 0,$
where the last inequality is due to $\sum_{s^{\prime}\in\mathcal{S}_{k,h}}p_{k,h}^{s^{\prime}}(\theta_{h}^{*})\log\frac{p_{k,h}^{s^{\prime}}(\theta_{h}^{*})}{p_{k,h}^{s^{\prime}}(\theta)}$ is the Kullback-Leibler divergence between $p_{k,h}(\theta_{h}^{*})$ and $p_{k,h}(\theta)$, which always is non-negative. ∎
B.2.2 Proof of Lemma 8
Proof.
For any positive semi-definite matrix $Z$,
$(x_{i}-x_{j})^{\top}Z(x_{i}-x_{j})=x_{i}^{\top}Zx_{i}-x_{i}^{\top}Zx_{j}-x_{j}^{\top}Zx_{i}+x_{j}^{\top}Zx_{j}\geq 0,\quad\forall x_{i},x_{j}\in\mathbb{R}^{d}.$ (23)
which implies $x_{i}^{\top}Zx_{i}+x_{j}^{\top}Zx_{j}\geq x_{i}^{\top}Zx_{j}+x_{j}^{\top}Zx_{i},\forall x_{i},x_{j}\in\mathbb{R}^{d}$. Let $x_{i}=(p_{k,h}^{s^{\prime}}(\theta)-y_{k,h}^{s^{\prime}})\phi_{k,h}^{s^{\prime}}$, we have
$\lVert g_{k,h}(\theta)\rVert_{Z}^{2}$
=\ = $\sum_{s^{\prime}\in\mathcal{S}_{k,h}}\sum_{s^{\prime\prime}\in\mathcal{S}_{k,h}}(p_{k,h}^{s^{\prime}}(\theta)-y_{k,h}^{s^{\prime}})(p_{k,h}^{s^{\prime\prime}}(\theta)-y_{k,h}^{s^{\prime\prime}})\phi_{k,h}^{s^{\prime}}Z\phi_{k,h}^{s^{\prime\prime}}$
=\ = $\sum_{s^{\prime}\in\mathcal{S}_{k,h}}(p_{k,h}^{s^{\prime}}(\theta)-y_{k,h}^{s^{\prime}})^{2}\phi_{k,h}^{s^{\prime}}Z\phi_{k,h}^{s^{\prime}}$
$\hskip 56.9055pt+\frac{1}{2}\sum_{s^{\prime}\in\mathcal{S}_{k,h}}\sum_{s^{\prime\prime}\in\mathcal{S}_{k,h}}(p_{k,h}^{s^{\prime}}(\theta)-y_{k,h}^{s^{\prime}})(p_{k,h}^{s^{\prime\prime}}(\theta)-y_{k,h}^{s^{\prime\prime}})(\phi_{k,h}^{s^{\prime}}Z\phi_{k,h}^{s^{\prime\prime}}+\phi_{k,h}^{s^{\prime\prime}}Z\phi_{k,h}^{s^{\prime}})$
[…]
=\ = $2\sum_{s^{\prime}\in\mathcal{S}_{k,h}}(p_{k,h}^{s^{\prime}}(\theta)-y_{k,h}^{s^{\prime}})^{2}\phi_{k,h}^{s^{\prime}}Z\phi_{k,h}^{s^{\prime}}$
[…]
This finishes the proof. ∎
B.2.3 Proof of Lemma 9
Proof.
Notice that $\left(\bar{g}_{i,h}(\theta_{i})-g_{i,h}(\theta_{i})\right)^{\top}(\theta_{i}-\theta^{*})$ is a martingale difference sequence. Also, we have
$\left|\left(\bar{g}_{i,h}(\theta_{i})-g_{i,h}(\theta_{i})\right)^{\top}(\theta_{i}-\theta^{*})\right|$ $\leq\left|\left(\bar{g}_{i,h}(\theta_{i})\right)^{\top}(\theta_{i}-\theta^{*})\right|+\left|\left(g_{i,h}(\theta_{i})\right)^{\top}(\theta_{i}-\theta^{*})\right|$
$\leq\lVert\bar{g}_{i,h}(\theta_{i})\rVert_{2}\lVert\theta_{i}-\theta^{*}\rVert_{2}+\lVert g_{i,h}(\theta_{i})\rVert_{2}\lVert\theta_{i}-\theta^{*}\rVert_{2}$
$\leq 4\sqrt{2}B,$
where the last inequality holds by $\lVert\bar{g}_{i,h}(\theta_{i})\rVert_{2}=\lVert\sum_{s^{\prime}\in\mathcal{S}_{i,h}}(p_{i,h}^{s^{\prime}}(\theta)-y_{i,h}^{s^{\prime}})\phi_{i,h}^{s^{\prime}}\rVert_{2}\leq\sqrt{2}$.
We define the martingale $M_{k,h}=\sum_{i=1}^{k}\left(\bar{g}_{i,h}(\theta_{i})-g_{i,h}(\theta_{i})\right)^{\top}(\theta_{i}-\theta^{*})$, and define the conditional variance $\sigma_{i}^{2}$ as
$\sigma_{k,h}^{2}$ $=\mathbb{E}[M_{k,h}^{2}\mid\mathcal{F}_{i-1}]$
$=\sum_{i=1}^{k}\mathbb{E}_{y_{k,h}}\left[\left(\bar{g}_{i,h}(\theta_{i})-g_{i,h}(\theta_{i})\right)^{\top}(\theta_{i}-\theta^{*})^{2}\right]$
$\leq\sum_{i=1}^{k}\mathbb{E}_{y_{k,h}}\left[\left(g_{i,h}(\theta_{i})\right)^{\top}(\theta_{i}-\theta^{*})^{2}\right]$
$\leq\sum_{i=1}^{k}\sum_{s^{\prime}\in\mathcal{S}_{i,h}}(\phi_{i,h}^{s^{\prime}})^{\top}(\theta_{i}-\theta^{*})^{2}=\sum_{i=1}^{k}\lVert\theta_{i}-\theta^{*}\rVert_{W_{i,h}}^{2}\triangleq A_{k,h}.$
where the first inequality is due to the fact that $\mathbb{E}[(\xi-\mathbb{E}[\xi])^{2}]\leq\mathbb{E}[\xi^{2}]$ for any random variable $\xi$. Note that $A_{k,h}$ is a random variable, so we cannot directly apply the Bernstein inequality to $M_{k,h}$. Instead, we consider the following two cases:(i) $A_{k,h}\leq\frac{4B^{2}}{kU}$ and (ii) $A_{k,h}>\frac{4B^{2}}{kU}$.
Case (i): $A_{k,h}=\sum_{i=1}^{k}\lVert\theta_{i}-\theta^{*}\rVert_{W_{i,h}}^{2}\leq\frac{4B^{2}}{kU}$. Then, we have
$M_{k,h}$ $=\sum_{i=1}^{k}\left(\bar{g}_{i,h}(\theta_{i})-g_{i,h}(\theta_{i})\right)^{\top}(\theta_{i}-\theta^{*})$
$=\sum_{i=1}^{k}\sum_{s^{\prime}\in\mathcal{S}_{i,h}}(p_{i,h}^{s^{\prime}}(\theta_{i})-y_{i,h}^{s^{\prime}})(\phi_{i,h}^{s^{\prime}})^{\top}(\theta_{i}-\theta^{*})$
$\leq\sum_{i=1}^{k}\sum_{s^{\prime}\in\mathcal{S}_{i,h}}\lvert(\phi_{i,h}^{s^{\prime}})^{\top}(\theta_{i}-\theta^{*})\rvert\leq\sqrt{kU\sum_{i=1}^{k}\sum_{s^{\prime}\in\mathcal{S}_{i,h}}((\phi_{i,h}^{s^{\prime}})^{\top}(\theta_{i}-\theta^{*}))^{2}}\leq 2B.$
Case (ii): $A_{k,h}=\sum_{i=1}^{k}\lVert\theta_{i}-\theta^{*}\rVert_{W_{j,h}}^{2}>\frac{4B^{2}}{kU}$. We have both a lower and upper bound for $A_{k,h}$, i.e., $\frac{4B^{2}}{kU}<A_{k,h}\leq 4B^{2}kU$. Then, we can use the peeling process to bound $M_{k,h}$ as follows:
$\Pr\left[M_{k,h}\geq 2\sqrt{\tau_{k,h}A_{k,h}}+\frac{8B\tau_{k,h}}{3}\right]$
=\ = $\Pr\left[M_{k,h}\geq 2\sqrt{\tau_{k,h}A_{k,h}}+\frac{8B\tau_{k,h}}{3},\frac{4B^{2}}{kU}<A_{k,h}\leq 4kUB^{2}\right]$
=\ = $\Pr\left[M_{k,h}\geq 2\sqrt{\tau_{k,h}A_{k,h}}+\frac{8B\tau_{k,h}}{3},\frac{4B^{2}}{kU}<A_{k,h}\leq 4kUB^{2},\sigma_{k,h}\leq A_{k,h}\right]$
[…]
where $m=2\log_{2}(kU)$, and the last inequality follows the Bernstein inequality for martingales. Combining above two cases, letting $\tau=\log{\frac{mk^{2}}{\delta/H}}$ and taking the union bound over $k$ and $h\in[H]$, we have with probability at least $1-\delta$, for any $k\in[K],h\in[H]$ it holds that
$M_{k,h}$ $=\sum_{i=1}^{k}\left(\bar{g}_{i,h}(\theta_{i})-g_{i,h}(\theta_{i})\right)^{\top}(\theta_{i}-\theta^{*})$
$\leq 2\sqrt{\tau_{k,h}A_{k,h}}+\frac{8B\tau_{k,h}}{3}+4\sqrt{2}B$
$\leq 2\sqrt{\sum_{i=1}^{k}\lVert\theta_{i}-\theta^{*}\rVert_{W_{i,h}}^{2}\log\frac{2k^{2}H\log(kU)}{\delta}}+8B\left(1+\log\frac{2k^{2}H\log(kU)}{\delta}\right)$
$=2\sqrt{\gamma_{k}\sum_{i=1}^{k}\lVert\theta_{i}-\theta^{*}\rVert_{W_{i,h}}^{2}}+8B\left(1+\gamma_{k}\right),$
where $\gamma_{k}=\log\frac{2k^{2}H\log(kU)}{\delta}$.
Then, applying $uv\leq cu^{2}+v^{2}/(4c)$ for any $c,u,v>0$ with $c=2/\kappa$, we have
$\sqrt{\gamma_{k}\sum_{i=1}^{k}\lVert\theta_{i}-\theta^{*}\rVert_{W_{i,h}}^{2}}\leq\frac{2\gamma_{k}}{\kappa}+\frac{\kappa}{8}\sum_{i=1}^{k}\lVert\theta_{i}-\theta^{*}\rVert_{W_{i,h}}^{2}$
Then we have
$\sum_{i=1}^{k}\left(\bar{g}_{i,h}(\theta_{i})-g_{i,h}(\theta_{i})\right)^{\top}(\theta_{i}-\theta^{*})\leq\frac{\kappa}{4}\sum_{i=1}^{k}\lVert\theta_{i}-\theta^{*}\rVert_{W_{i,h}}^{2}+(\frac{4}{\kappa}+8B)\gamma_{k}.$
This finishes the proof. ∎
Appendix C Proof of Lemma 2
Proof.
The gradient of $p_{s,a}(\theta)$ is given by $\nabla p_{s,a}^{s^{\prime}}(\theta)=p_{s,a}^{s^{\prime}}(\theta)\phi_{s,a}^{s^{\prime}}-p_{s,a}^{s^{\prime}}(\theta)\sum_{s^{\prime\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime\prime}}(\theta)\phi_{s,a}^{s^{\prime\prime}}$. By the mean value theorem, there exists $\bar{\theta}=\nu\theta_{h}^{*}+(1-\nu)\widehat{\theta}_{k,h}$ for some $\nu\in[0,1]$, such that
$\bigg{|}{\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\widehat{\theta}_{k,h})V(s^{\prime})-\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\theta_{h}^{*})V(s^{\prime})}\bigg{|}$
=\ = $\bigg{|}\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\bar{\theta})\phi_{s,a}^{s^{\prime}}(\widehat{\theta}_{k,h}-\theta_{h}^{*})V(s^{\prime})-\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\bar{\theta})\sum_{s^{\prime\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime\prime}}(\bar{\theta})\phi_{s,a}^{s^{\prime\prime}}(\widehat{\theta}_{k,h}-\theta_{h}^{*})V(s^{\prime})\bigg{|}$
=\ = $\bigg{|}\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\bar{\theta})\phi_{s,a}^{s^{\prime}}(\widehat{\theta}_{k,h}-\theta_{h}^{*})\Big{(}V(s^{\prime})-\sum_{s^{\prime\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime\prime}}(\bar{\theta})V(s^{\prime\prime})\Big{)}\bigg{|}$
[…]
where the first inequality is by $V:\mathcal{S}\to[0,H]$, the second inequality is by the definition of $\phi_{k,h}^{s^{\prime}}$, the third inequality is by the Holder’s inequality, the last inequality is by Lemma 1. ∎
Appendix D Proof of Theorem 1
D.1 Main Proof
To prove the theorem, we first introduce the following lemma.
Lemma 10.
Suppose for all $(h,s,a)\in[H]\times\mathcal{S}\times\mathcal{A}$, it holds that
$\bigg{|}{\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\widehat{\theta}_{k,h})V(s^{\prime})-\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\theta_{h}^{*})V(s^{\prime})}\bigg{|}\leq\Gamma_{h,s,a}.$ (25)
Define
$\widehat{Q}_{k,h}(s,a)=\bigg{[}r_{h}(s,a)+\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\widehat{\theta}_{k,h})\widehat{V}_{k,h+1}(s^{\prime})+\Gamma_{h,s,a}\bigg{]}_{[0,H]}.$ (26)
Then, for any $\delta\in(0,1]$, with probability at least $1-\delta$, it holds that
$\operatorname*{Reg}(K)\leq 2\sum_{k=1}^{K}\sum_{h=1}^{H}\Gamma_{h,s_{k,h},a_{k,h}}+H\sqrt{2KH\log(2/\delta)}.$
Proof of Theorem 1.
Substituting $\Gamma_{h,s,a}=H\widehat{\beta}_{k}\max_{s^{\prime}\in\mathcal{S}_{h,s,a}}\lVert\phi_{s,a}^{s^{\prime}}\rVert_{\widehat{\Sigma}_{k,h}^{-1}}$ into Lemma 10, we have
$\operatorname*{Reg}(K)$ $\leq 2H\widehat{\beta}_{k}\sum_{k=1}^{K}\sum_{h=1}^{H}\max_{s^{\prime}\in\mathcal{S}_{k,h}}\lVert\phi_{k,h}^{s^{\prime}}\rVert_{\widehat{\Sigma}_{k,h}^{-1}}+H\sqrt{2KH\log(2/\delta)}$
$\leq 2H\widehat{\beta}_{k}\sum_{h=1}^{H}\sqrt{K\sum_{k=1}^{K}\max_{s^{\prime}\in\mathcal{S}_{k,h}}\lVert\phi_{k,h}^{s^{\prime}}\rVert_{\widehat{\Sigma}_{k,h}^{-1}}^{2}}+H\sqrt{2KH\log(2/\delta)}$
$\leq 2H^{2}\widehat{\beta}_{k}\sqrt{\frac{4dK}{\kappa}\log\left(1+\frac{KU}{d\lambda}\right)}+H\sqrt{2KH\log(2/\delta)}\leq\widetilde{\mathcal{O}}({\kappa}^{-1}dH^{2}\sqrt{K}),$
where the second inequality holds by the Cauchy-Schwarz inequality and the third inequality holds by Lemma 6. This finishes the proof. ∎
D.2 Proof of Auxiliary Lemmas
In this section, we provide the proofs of the lemmas used in Appendix D.1.
First, we introduce the following lemma.
Lemma 11.
Suppose (25) and (26) in Lemma 10 holds for all $k\in[K],h\in[H]$. Then, for any $(s,a,h)\in\mathcal{S}\times\mathcal{A}\times[H]$, it holds that
$Q_{h}^{*}(s,a)\leq\widehat{Q}_{k,h}(s,a)\leq r_{h}(s,a)+\mathbb{P}_{h}V_{k,h+1}(s,a)+2\Gamma_{h,s,a}.$
Proof.
First, we prove the left-hand side of the lemma. We prove this by backward induction on $h$. For the stage $h=H$, by definition, we have
$\widehat{Q}_{k,H}(s,a)=r_{H}(s,a)=Q_{H}^{*}(s,a),\quad\widehat{V}_{k,H+1}(s)=0=V_{H+1}^{*}(s).$
Suppose the statement holds for $h+1$, we show it holds for $h$. By definition, if $\widehat{Q}_{k,h}(s,a)=H$, this holds trivially. Otherwise, we have
$\widehat{Q}_{k,h}(s,a)$ $=r_{h}(s,a)+p_{s,a}(\widehat{\theta}_{k,h})\widehat{V}_{k,h+1}+\Gamma_{h,s,a}$
$\geq r_{h}(s,a)+p_{s,a}(\widehat{\theta}_{k,h})V_{k}^{*}+\Gamma_{h,s,a}\geq r_{h}(s,a)+p_{s,a}(\theta_{h}^{*})V_{k}^{*}=Q_{h}^{*}(s,a).$
where the first inequality is by the induction hypothesis, and the second inequality is by (25).
Then, we prove the right-hand side of the lemma. By the definition of $\widehat{Q}_{k,h}(s,a)$, we have
$\widehat{Q}_{k,h}(s,a)=r_{h}(s,a)+p_{s,a}(\widehat{\theta}_{k,h})\widehat{V}_{k,h+1}+\Gamma_{h,s,a}\leq r_{h}(s,a)+p_{s,a}(\theta_{h}^{*})\widehat{V}_{k,h+1}+2\Gamma_{h,s,a},$
where the inequality is by (25). ∎
D.2.1 Proof of Lemma 10
Proof.
By the definition of $\operatorname*{Reg}(K)=\sum_{k=1}^{K}V_{1}^{*}(s_{k,1})-\sum_{k=1}^{K}V_{1}^{\pi_{k}}(s_{k,1})$, we have
$\sum_{k=1}^{K}V_{1}^{*}(s_{k,1})-\sum_{k=1}^{K}V_{1}^{\pi_{k}}(s_{k,1})=$ $\sum_{k=1}^{K}Q_{1}^{*}(s_{k,1},\pi^{*}(s_{k,1}))-\sum_{k=1}^{K}V_{1}^{\pi_{k}}(s_{k,1})$
[…]
where the first inequality is by Lemma 11 and the second is by $a_{k,h}=\operatorname*{arg\,max}_{a\in\mathcal{A}}\widehat{Q}_{k,h}(s_{k,h},a)$.
By the right-hand side of Lemma 11, we have
$\widehat{Q}_{1}(s_{k,1},a_{k,1})-V_{1}^{\pi_{k}}(s_{k,1})$
=\ = $r(s_{k,1},a_{k,1})+\mathbb{P}_{1}\widehat{V}_{k,2}(s_{k,1},a_{k,1})++2\Gamma_{h,s_{k,1},a_{k,1}}-r(s_{k,1},a_{k,1})-\mathbb{P}_{1}V_{2}^{\pi_{k}}(s_{k,1},a_{k,1})$
[…]
Define $\mathcal{M}_{k,h}=\mathbb{P}_{h}(\widehat{V}_{k,h+1}-V_{h+1}^{\pi_{k}})(s_{k,h},a_{k,h})-(\widehat{V}_{k,h+1}-V_{h+1}^{\pi_{k}})(s_{k,h+1})$. Applying this recursively, we have
$\widehat{Q}_{1}(s_{k,1},a_{k,1})-V_{1}^{\pi_{k}}(s_{k,1})\leq 2\sum_{h=1}^{H}\Gamma_{h,s_{k,h},a_{k,h}}+\sum_{h=1}^{H}\mathcal{M}_{k,h}$
Summing over $k$, we have for any $\delta\in(0,1]$, with probability at least $1-\delta$, it holds that
$\operatorname*{Reg}(K)\leq 2\sum_{k=1}^{K}\sum_{h=1}^{H}\Gamma_{h,s_{k,h},a_{k,h}}+\sum_{k=1}^{K}\sum_{h=1}^{H}\mathcal{M}_{k,h}\leq 2\sum_{k=1}^{K}\sum_{h=1}^{H}\Gamma_{h,s_{k,h},a_{k,h}}+H\sqrt{2KH\log(2/\delta)}$
where the inequality holds by the Azuma-Hoeffding inequality as $\mathcal{M}_{k,h}$ is a martingale difference sequence with $\mathcal{M}_{k,h}\leq 2H$. This finishes the proof. ∎
Appendix E Proof of Lemma 3
Proof.
The proof is similar to the proof of Theorem 3 in Zhang and Sugiyama [2023] and Lemma 1 of Lee and Oh [2024]. Define
$\widetilde{f}_{k,h}(\theta)=f_{k,h}(\theta_{k,h})+\langle g_{k,h}(\widetilde{\theta}_{k,h}),\theta-\widetilde{\theta}_{k,h}\rangle+\frac{1}{2}\lVert\theta-\widetilde{\theta}_{k,h}\rVert_{H_{k,h}(\widetilde{\theta}_{k,h})},$
which is a second order approximation of the original function $f_{k,h}(\theta)$ at $\widetilde{\theta}_{k,h}$. Then, the update rule in (8) can be rewritten as
$\widetilde{\theta}_{k+1,h}=\operatorname*{arg\,min}_{\theta\in\Theta}\widetilde{f}_{k,h}(\theta)+\frac{1}{2\eta}\lVert\theta-\widetilde{\theta}_{k,h}\rVert_{\mathcal{H}_{k,h}}.$ (27)
Then, we present the following lemma.
Lemma 12 (Lemma E.1 of Lee and Oh [2024]).
For the update rule defined in (27) as
$\widetilde{\theta}_{k+1,h}=\operatorname*{arg\,min}_{\theta\in\Theta}\widetilde{f}_{k,h}(\theta)+\frac{1}{2\eta}\lVert\theta-\widetilde{\theta}_{k,h}\rVert_{\mathcal{H}_{k,h}},$
let $\eta=\frac{1}{2}\log(U+1)+(B+1)$, we have
$\lVert\widetilde{\theta}_{k+1,h}-\theta_{h}^{*}\rVert_{\mathcal{H}_{k+1,h}}^{2}\leq$ $2\eta\left(\sum_{i=1}^{k}f_{i,h}(\theta_{h}^{*})-\sum_{i=1}^{k}f_{i,h}(\widetilde{\theta}_{i+1,h})\right)+4\lambda B$
$+12\sqrt{2}B\eta\sum_{i=1}^{k}\lVert\widetilde{\theta}_{i+1,h}-\widetilde{\theta}_{i,h}\rVert_{2}^{2}-\sum_{i=1}^{k}\lVert\widetilde{\theta}_{i+1,h}-\widetilde{\theta}_{i,h}\rVert_{\mathcal{H}_{i,h}}^{2}.$
Then, we bound the right-hand side of the above lemma separately. First, we bound the first term. Define the softmax function as follows.
$[\sigma_{k,h}(z)]_{s}=\frac{\exp([z])_{s}}{1+\sum_{s\in\dot{\mathcal{S}}_{k,h}}\exp([z])_{s}},\quad\forall s\in\mathcal{S}_{k,h}.$
Then, the loss function in (2) can be rewritten as
$f_{k,h}(z_{k,h},y_{k,h})=\sum_{s^{\prime}\in\mathcal{S}_{k,h}}\operatorname{\mathds{1}}[y_{k,h}^{s^{\prime}}=1]\log\left(\frac{1}{[\sigma_{k,h}(z_{k,h})]_{s^{\prime}}}\right).$
Define a pseudo-inverse function of $\sigma_{k,h}(\cdot)$ as
$[\sigma_{k,h}^{-1}(p)]_{s^{\prime}}=\log\left(\frac{[p]_{s^{\prime}}}{1-\lVert p\rVert_{1}}\right),\forall p\in\{p\in[0,1]^{N_{k,h}}\mid\lVert p\rVert_{1}<1\}.$
Then, we decompose the first term as follows.
$\sum_{i=1}^{k}f_{i,h}(\theta_{h}^{*})-\sum_{i=1}^{k}f_{i,h}(\widetilde{\theta}_{i+1,h})=\underbrace{\sum_{i=1}^{k}f_{i,h}(\theta_{h}^{*})-\sum_{i=1}^{k}f_{i,h}(z_{i,h},y_{i,h})}_{(a)}+\underbrace{\sum_{i=1}^{k}f_{i,h}(z_{i,h},y_{i,h})-\sum_{i=1}^{k}f_{i,h}(\widetilde{\theta}_{i+1,h})}_{(b)}$
where $z_{k,h}=\sigma^{-1}_{k,h}(\mathbb{E}_{\theta\sim P_{k,h}}[\sigma_{k,h}((\phi_{k,h}^{s^{\prime}})^{\top}\theta)_{s^{\prime}\in\dot{\mathcal{S}}_{k,h}}])$, $P_{k,h}\triangleq\mathcal{N}(\widetilde{\theta}_{k,h},(1+c\mathcal{H}_{k,h}^{-1}))$ is the Gaussian distribution with mean $\widetilde{\theta}_{k,h}$ and covariance $(1+c\mathcal{H}_{k,h}^{-1})$ where $c$ is a constant to be specified later.
First, we bound the term $(a)$ as follows.
Lemma 13 (Lemma E.2 of Lee and Oh [2024]).
Let $\delta\in(0,1]$ and $\lambda\geq 1$, for all $k\in[K],h\in[H]$, with probability at least $1-\delta$, we have
$\sum_{i=1}^{k}f_{i,h}(\theta_{h}^{*})-\sum_{i=1}^{k}f_{i,h}(z_{i,h},y_{i,h})$
[…]
Then, we bound the term $(b)$ as follows.
Lemma 14 (Lemma E.3 of Lee and Oh [2024]).
For any $c\geq 0$, let $\lambda\geq\max\{2,72cd\}$, then for all $k\in[K],h\in[H]$, we have
$\sum_{i=1}^{k}f_{i,h}(z_{i,h},y_{i,h})-\sum_{i=1}^{k}f_{i,h}(\widetilde{\theta}_{i+1,h})\leq\frac{1}{2c}\sum_{i=1}^{k}\left\|\widetilde{\theta}_{i,h}-\theta_{i+1,h}\right\|_{\mathcal{H}_{i,h}}^{2}+\sqrt{6}cd\log\left(1+\frac{k+1}{2\lambda}\right).$
Now, we are ready to prove Lemma 3. Combining Lemma 12, Lemma 13, and Lemma 14, we have
$\lVert\widetilde{\theta}_{k+1,h}-\theta_{h}^{*}\rVert_{\mathcal{H}_{k+1,h}}^{2}$
[…]
$+$ $2+\sqrt{6}cd\log\left(1+\frac{k+1}{2\lambda}\right)\Bigg{]}+4\lambda B+12\sqrt{2}B\eta\sum_{i=1}^{k}\lVert\widetilde{\theta}_{i+1,h}-\widetilde{\theta}_{i,h}\rVert_{2}^{2}+(\frac{\eta}{c}-1)\sum_{i=1}^{k}\lVert\widetilde{\theta}_{i+1,h}+\widetilde{\theta}_{i,h}\rVert_{\mathcal{H}_{i,h}}^{2}$
[…]
$+2+\sqrt{6}cd\log\left(1+\frac{k+1}{2\lambda}\right)\Bigg{]}+4\lambda B,$
where the second inequality holds by setting $c={7\eta}/{6}$ and $\lambda\geq\max\{84\sqrt{2}\eta B,84d\eta\}$, we have
$12\sqrt{2}B\eta\sum_{i=1}^{k}\lVert\widetilde{\theta}_{i+1,h}-\widetilde{\theta}_{i,h}\rVert_{2}^{2}+\Big{(}\frac{\eta}{c}-1\Big{)}\sum_{i=1}^{k}\lVert\widetilde{\theta}_{i+1,h}+\widetilde{\theta}_{i,h}\rVert_{\mathcal{H}_{i,h}}^{2}$
=\ = $12\sqrt{2}B\eta\sum_{i=1}^{k}\lVert\widetilde{\theta}_{i+1,h}-\widetilde{\theta}_{i,h}\rVert_{2}^{2}-\frac{1}{7}\sum_{i=1}^{k}\lVert\widetilde{\theta}_{i+1,h}+\widetilde{\theta}_{i,h}\rVert_{\mathcal{H}_{i,h}}^{2}$
[…]
Thus, by setting $\eta=\frac{1}{2}\log(U+1)+(B+1)$, $\lambda=84\sqrt{2}\eta(B+d)$, we have $\lVert\widetilde{\theta}_{k+1,h}-\theta_{h}^{*}\rVert_{\mathcal{H}_{k+1,h}}\leq\mathcal{O}\big{(}\sqrt{d}\log U\log(KH/\delta)\big{)}\triangleq\widetilde{\beta}_{k}$. This finishes the proof. ∎
Appendix F Proof of Theorem 2
Proof.
By the same analysis as the proof of Lemma 2 in Appendix C, we have
$\bigg{|}{\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\widetilde{\theta}_{k,h})V(s^{\prime})-\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\theta_{h}^{*})V(s^{\prime})}\bigg{|}\leq$ $H\max_{s^{\prime}\in\mathcal{S}_{h,s,a}}\phi_{s,a}^{s^{\prime}}(\widetilde{\theta}_{k,h}-\theta_{h}^{*})$
[…]
where the first inequality holds by (24), the second inequality is by Holder’s inequality, and the third inequality is by Lemma 3.
Then, substituting $\Gamma_{h,s,a}=H\widetilde{\beta}_{k}\max_{s^{\prime}\in\mathcal{S}_{h,s,a}}\lVert\phi_{s,a}^{s^{\prime}}\rVert_{\mathcal{H}_{k,h}^{-1}}$ into Lemma 10, we have
$\sum_{k=1}^{K}V_{1}^{*}(s_{k,1})-\sum_{k=1}^{K}V_{1}^{\pi_{k}}(s_{k,1})\leq$ $2H\widetilde{\beta}_{k}\sum_{k=1}^{K}\sum_{h=1}^{H}\max_{s^{\prime}\in\mathcal{S}_{k,h}}\lVert\phi_{k,h}^{s^{\prime}}\rVert_{\mathcal{H}_{k,h}^{-1}}+H\sqrt{2KH\log(2/\delta)}$
[…]
where the second inequality holds by the Cauchy-Schwarz inequality, and the third inequality is by Lemma 6. This finishes the proof. ∎
Appendix G Proof of Lemma 4
Proof.
By the second-order Taylor expansion, there exists $\bar{\theta}=\nu\theta_{h}^{*}+(1-\nu)\widetilde{\theta}_{k,h}$ for some $\nu\in[0,1]$, such that
${\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\widetilde{\theta}_{k,h})V(s^{\prime})-\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\theta_{h}^{*})V(s^{\prime})}$
=\ = $\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}\nabla p_{s,a}^{s^{\prime}}(\widetilde{\theta}_{k,h})^{\top}(\theta_{h}^{*}-\widetilde{\theta}_{k,h})V(s^{\prime})+\frac{1}{2}\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}(\widehat{\theta}_{k,h}-\theta_{h}^{*})^{\top}\nabla^{2}p_{s,a}^{s^{\prime}}(\bar{\theta})(\widehat{\theta}_{k,h}-\theta_{h}^{*})V(s^{\prime})$
The gradient of $p_{s,a}(\theta)$ is given by $\nabla p_{s,a}^{s^{\prime}}(\theta)=p_{s,a}^{s^{\prime}}(\theta)\phi_{s,a}^{s^{\prime}}-p_{s,a}^{s^{\prime}}(\theta)\sum_{s^{\prime\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime\prime}}(\theta)\phi_{s,a}^{s^{\prime\prime}}$. For the first-order term, we have
$\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}\nabla p_{s,a}^{s^{\prime}}(\widetilde{\theta}_{k,h})^{\top}(\theta_{h}^{*}-\widetilde{\theta}_{k,h})V(s^{\prime})$
=\ = $\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\widetilde{\theta}_{k,h})\phi_{s,a}^{s^{\prime}}(\theta_{h}^{*}-\widetilde{\theta}_{k,h})V(s^{\prime})-\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\widetilde{\theta}_{k,h})\sum_{s^{\prime\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime\prime}}(\widetilde{\theta}_{k,h})\phi_{s,a}^{s^{\prime\prime}}(\theta_{h}^{*}-\widetilde{\theta}_{k,h})V(s^{\prime})$
[…]
where in the first inequality, we denote $\mathcal{S}_{h,s,a}^{+}$ as the subset of $\mathcal{S}_{h,s,a}$ such that $\phi_{s,a}^{s^{\prime}}(\theta_{h}^{*}-\widetilde{\theta}_{k,h})-\sum_{s^{\prime\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime\prime}}(\widetilde{\theta}_{k,h})\phi_{s,a}^{s^{\prime\prime}}(\theta_{h}^{*}-\widetilde{\theta}_{k,h})$ is non-negative, the second inequality holds by the Holder’s inequality, and the third inequality is by the confidence set in Lemma 3.
For the second-order term, we define $u_{s,a}^{s^{\prime}}(\theta)=\phi_{s,a}^{s^{\prime}}\theta$, and $p_{s,a}^{s^{\prime}}(u)=\frac{\exp(u_{s,a}^{s^{\prime}})}{1+\sum_{s^{\prime\prime}}\exp(u_{s,a}^{s^{\prime\prime}})}$, further define
$F(u)=\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}\frac{\exp(u_{s,a}^{s^{\prime}})}{1+\sum_{s^{\prime\prime}\in\mathcal{S}_{h,s,a}}\exp(u_{s,a}^{s^{\prime\prime}})},\widetilde{F}(u)=\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}\frac{\exp(u_{s,a}^{s^{\prime}})V(s^{\prime})}{1+\sum_{s^{\prime\prime}\in\mathcal{S}_{h,s,a}}\exp(u_{s,a}^{s^{\prime\prime}})}.$
Then, we have
$\frac{1}{2}\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}(\widehat{\theta}_{k,h}-\theta_{h}^{*})^{\top}\nabla^{2}p_{s,a}^{s^{\prime}}(\bar{\theta})(\widehat{\theta}_{k,h}-\theta_{h}^{*})V(s^{\prime})$
=\ = $\frac{1}{2}\big{(}u(\widehat{\theta}_{k,h})-u(\theta_{h}^{*})\big{)}^{\top}\nabla^{2}\widetilde{F}(u(\bar{\theta}))\big{(}u(\widehat{\theta}_{k,h})-u(\theta_{h}^{*})\big{)}$
=\ = $\frac{1}{2}\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}\sum_{s^{\prime\prime}\in\mathcal{S}_{h,s,a}}\big{(}u_{s,a}^{s^{\prime}}(\widehat{\theta}_{k,h})-u_{s,a}^{s^{\prime}}(\theta_{h}^{*})\big{)}^{\top}\frac{\partial^{2}\widetilde{F}(u(\bar{\theta}))}{\partial s^{\prime}\partial s^{\prime\prime}}\big{(}u_{s,a}^{s^{\prime\prime}}(\widehat{\theta}_{k,h})-u_{s,a}^{s^{\prime\prime}}(\theta_{h}^{*})\big{)}$
[…]
where the inequality holds by $V(s)\in[0,H],\forall s$. According to Lemma 18, we have (omit the subscript $\mathcal{S}_{h,s,a}$ for simplicity):
$\frac{H}{2}\sum_{s^{\prime}}\sum_{s^{\prime\prime}}\big{\lvert}u_{s,a}^{s^{\prime}}(\widehat{\theta}_{k,h})-u_{s,a}^{s^{\prime}}(\theta_{h}^{*})\big{\rvert}\frac{\partial^{2}F(u(\bar{\theta}))}{\partial s^{\prime}\partial s^{\prime\prime}}\big{\lvert}u_{s,a}^{s^{\prime\prime}}(\widehat{\theta}_{k,h})-u_{s,a}^{s^{\prime\prime}}(\theta_{h}^{*})\big{\rvert}$
[…]
$+\frac{3H}{2}\sum_{s^{\prime}}\big{(}u_{s,a}^{s^{\prime}}(\widehat{\theta}_{k,h})-u_{s,a}^{s^{\prime}}(\theta_{h}^{*})\big{)}^{2}p_{s,a}^{s^{\prime}}(u(\bar{\theta})).$ (29)
To bound the first term, by applying the AM-GM inequality, we obtain
$H\sum_{s^{\prime}}\sum_{s^{\prime\prime}\neq s^{\prime}}\big{\lvert}u_{s,a}^{s^{\prime}}(\widehat{\theta}_{k,h})-u_{s,a}^{s^{\prime}}(\theta_{h}^{*})\big{\rvert}p_{s,a}^{s^{\prime}}\big{(}u(\bar{\theta})\big{)}p_{s,a}^{s^{\prime\prime}}\big{(}u(\bar{\theta})\big{)}\big{\lvert}u_{s,a}^{s^{\prime\prime}}(\widehat{\theta}_{k,h})-u_{s,a}^{s^{\prime\prime}}(\theta_{h}^{*})\big{\rvert}$
[…]
$\hskip 56.9055pt+\frac{H}{2}\sum_{s^{\prime}}\sum_{s^{\prime\prime}}\big{(}u_{s,a}^{s^{\prime\prime}}(\widehat{\theta}_{k,h})-u_{s,a}^{s^{\prime\prime}}(\theta_{h}^{*})\big{)}^{2}p_{s,a}^{s^{\prime}}\big{(}u(\bar{\theta})\big{)}p_{s,a}^{s^{\prime\prime}}\big{(}u(\bar{\theta})\big{)}$
[…]
Plugging (30) into (29), we have
$\frac{H}{2}\sum_{s^{\prime}}\sum_{s^{\prime\prime}}\big{\lvert}u_{s,a}^{s^{\prime}}(\widehat{\theta}_{k,h})-u_{s,a}^{s^{\prime}}(\theta_{h}^{*})\big{\rvert}\frac{\partial^{2}F(u(\bar{\theta}))}{\partial s^{\prime}\partial s^{\prime\prime}}\big{\lvert}u_{s,a}^{s^{\prime\prime}}(\widehat{\theta}_{k,h})-u_{s,a}^{s^{\prime\prime}}(\theta_{h}^{*})\big{\rvert}$
[…]
=\ = $\frac{5H}{2}\sum_{s^{\prime}}\big{(}(\phi_{s,a}^{s^{\prime}})^{\top}(\widehat{\theta}_{k,h}-\theta_{h}^{*})\big{)}^{2}p_{s,a}^{s^{\prime}}(u(\bar{\theta}))$
[…]
where the last inequality holds by Lemma 3.
Combining (28) and (31) finishes the proof. ∎
Appendix H Proof of Theorem 3
Proof.
Combining Lemma 4 and Lemma 10, we have
$\sum_{k=1}^{K}V_{1}^{*}(s_{k,1})-\sum_{k=1}^{K}V_{1}^{\pi_{k}}(s_{k,1})$ $\leq 2\sum_{k=1}^{K}\sum_{h=1}^{H}(\epsilon_{k,h}^{\mathtt{fst}}+\epsilon_{k,h}^{\mathtt{snd}})+H\sqrt{2KH\log(2/\delta)}$
where
$\epsilon_{k,h}^{\mathtt{fst}}=H\widetilde{\beta}_{k}\sum_{s^{\prime}\in\mathcal{S}_{k,h}}p_{k,h}^{s^{\prime}}(\widetilde{\theta}_{k,h})\Big{\lVert}\phi_{k,h}^{s^{\prime}}-\sum_{s^{\prime\prime}\in\mathcal{S}_{k,h}}p_{k,h}^{s^{\prime\prime}}(\widetilde{\theta}_{k,h})\phi_{k,h}^{s^{\prime\prime}}\Big{\rVert}_{\mathcal{H}_{k,h}^{-1}},\epsilon_{k,h}^{\mathtt{snd}}=\frac{5}{2}H\widetilde{\beta}_{k}^{2}\max_{s^{\prime}\in\mathcal{S}_{k,h}}\lVert\phi_{k,h}^{s^{\prime}}\rVert_{\mathcal{H}_{k,h}^{-1}}^{2}.$
Next, we bound $\epsilon_{k,h}^{\mathtt{fst}}$ and $\epsilon_{k,h}^{\mathtt{snd}}$ respectively.
Bounding $\epsilon_{k,h}^{\mathtt{fst}}$.
For simplicity, we denote
$\mathbb{E}_{\theta}[\phi_{k,h}^{s^{\prime}}]=\mathbb{E}_{s^{\prime}\sim p_{k,h}^{s}(\theta)}[\phi_{k,h}^{s^{\prime}}],\quad\bar{\phi}_{s,a}^{s^{\prime}}=\phi_{s,a}^{s^{\prime}}-\mathbb{E}_{\widetilde{\theta}_{k,h}}[\phi_{k,h}^{s^{\prime}}],\quad\widetilde{\phi}_{s,a}^{s^{\prime}}=\phi_{s,a}^{s^{\prime}}-\mathbb{E}_{\widetilde{\theta}_{k+1,h}}[\phi_{k,h}^{s^{\prime}}]$
Then, we have
$\sum_{s^{\prime}\in\mathcal{S}_{k,h}}p_{k,h}^{s^{\prime}}(\widetilde{\theta}_{k,h})\Big{\lVert}\phi_{k,h}^{s^{\prime}}-\sum_{s^{\prime\prime}\in\mathcal{S}_{k,h}}p_{k,h}^{s^{\prime\prime}}(\widetilde{\theta}_{k,h})\phi_{k,h}^{s^{\prime\prime}}\Big{\rVert}_{\mathcal{H}_{k,h}^{-1}}=\sum_{s^{\prime}\in\mathcal{S}_{k,h}}p_{k,h}^{s^{\prime}}(\widetilde{\theta}_{k,h})\big{\lVert}\bar{\phi}_{k,h}^{s^{\prime}}\big{\rVert}_{\mathcal{H}_{k,h}^{-1}}$
[…]
=\ = $\underbrace{\sum_{s^{\prime}\in\mathcal{S}_{k,h}}p_{k,h}^{s^{\prime}}(\widetilde{\theta}_{k,h})\big{\lVert}\bar{\phi}_{k,h}^{s^{\prime}}-\widetilde{\phi}_{k,h}^{s^{\prime}}\big{\rVert}_{\mathcal{H}_{k,h}^{-1}}}_{(c)}+\underbrace{\sum_{s^{\prime}\in\mathcal{S}_{k,h}}(p_{k,h}^{s^{\prime}}(\widetilde{\theta}_{k,h})-p_{k,h}^{s^{\prime}}(\widetilde{\theta}_{k+1,h}))\big{\lVert}\widetilde{\phi}_{k,h}^{s^{\prime}}\big{\rVert}_{\mathcal{H}_{k,h}^{-1}}}_{(d)}$
$+\underbrace{\sum_{s^{\prime}\in\mathcal{S}_{k,h}}p_{k,h}^{s^{\prime}}(\widetilde{\theta}_{k+1,h})\big{\lVert}\widetilde{\phi}_{k,h}^{s^{\prime}}\big{\rVert}_{\mathcal{H}_{k,h}^{-1}}}_{(e)}$
Next, we bound these terms separately.
For the first term $(c)$, we have
$\big{\lVert}\bar{\phi}_{k,h}^{s^{\prime}}-\widetilde{\phi}_{k,h}^{s^{\prime}}\big{\rVert}_{\mathcal{H}_{k,h}^{-1}}$
=\ = $\Big{\lVert}\sum_{s^{\prime\prime}\in\mathcal{S}_{k,h}}\left(p_{k,h}^{s^{\prime\prime}}(\widetilde{\theta}_{k+1,h})-p_{k,h}^{s^{\prime\prime}}(\widetilde{\theta}_{k,h})\right)\phi_{k,h}^{s^{\prime\prime}}\Big{\rVert}_{\mathcal{H}_{k,h}^{-1}}$
=\ = $\Big{\lVert}\sum_{s^{\prime\prime}\in\mathcal{S}_{k,h}}\left(\nabla p_{k,h}^{s^{\prime\prime}}(\xi_{k,h})^{\top}(\widetilde{\theta}_{k+1,h}-\widetilde{\theta}_{k,h})\right)\phi_{k,h}^{s^{\prime\prime}}\Big{\rVert}_{\mathcal{H}_{k,h}^{-1}}$
[…]
=\ = $\sum_{s^{\prime\prime}\in\mathcal{S}_{k,h}}\Big{\lvert}\Big{(}p_{k,h}^{s^{\prime\prime}}(\xi_{k,h})\phi_{k,h}^{s^{\prime\prime}}-p_{k,h}^{s^{\prime\prime}}(\xi_{k,h})\sum_{s^{\prime\prime\prime}\in\mathcal{S}_{k,h}}p_{k,h}^{s^{\prime\prime\prime}}(\xi_{k,h})\phi_{k,h}^{s^{\prime\prime\prime}}\Big{)}^{\top}(\widetilde{\theta}_{k+1,h}-\widetilde{\theta}_{k,h})\Big{\rvert}\big{\lVert}\phi_{k,h}^{s^{\prime\prime}}\big{\rVert}_{\mathcal{H}_{k,h}^{-1}}$
[…]
$\hskip 28.45274pt+\sum_{s^{\prime\prime}\in\mathcal{S}_{k,h}}p_{k,h}^{s^{\prime\prime}}(\xi_{k,h})\big{\lVert}\phi_{k,h}^{s^{\prime\prime}}\big{\rVert}_{\mathcal{H}_{k,h}^{-1}}\sum_{s^{\prime\prime\prime}\in\mathcal{S}_{k,h}}p_{k,h}^{s^{\prime\prime\prime}}(\xi_{k,h})\big{\lvert}(\phi_{k,h}^{s^{\prime\prime\prime}})^{\top}(\widetilde{\theta}_{k+1,h}-\widetilde{\theta}_{k,h})\big{\rvert}$
[…]
$\hskip 28.45274pt+\sum_{s^{\prime\prime}\in\mathcal{S}_{k,h}}p_{k,h}^{s^{\prime\prime}}(\xi_{k,h})\big{\lVert}\phi_{k,h}^{s^{\prime\prime}}\big{\rVert}_{\mathcal{H}_{k,h}^{-1}}\sum_{s^{\prime\prime\prime}\in\mathcal{S}_{k,h}}p_{k,h}^{s^{\prime\prime\prime}}(\xi_{k,h})\big{\lVert}\phi_{k,h}^{s^{\prime\prime\prime}}\big{\rVert}_{\mathcal{H}_{k,h}^{-1}}\big{\lVert}\widetilde{\theta}_{k+1,h}-\widetilde{\theta}_{k,h}\big{\rVert}_{\mathcal{H}_{k,h}}$
[…]
where the fifth inequality is by Cauchy-Schwarz inequality and the fourth inequality is because by Lemma 17, since ~$\widetilde{\mathcal{H}}_{k,h}\succeq\mathcal{H}_{k,h}\succeq\lambda I_{d}$, we have
$\big{\lVert}\widetilde{\theta}_{k+1,h}-\widetilde{\theta}_{k,h}\big{\rVert}_{\mathcal{H}_{k,h}}\leq\big{\lVert}\widetilde{\theta}_{k+1,h}-\widetilde{\theta}_{k,h}\big{\rVert}_{\widetilde{\mathcal{H}}_{k,h}}\leq 2\eta\lVert g_{k,h}(\widetilde{\theta}_{k,h})\rVert_{\widetilde{\mathcal{H}}_{k,h}^{-1}}\leq\frac{2\eta}{\sqrt{\lambda}}\lVert g_{k,h}(\widetilde{\theta}_{k,h})\rVert_{2},$
and since $g_{k,h}(\theta)=\sum_{s^{\prime}\in\mathcal{S}_{k,h}}(p_{k,h}^{s^{\prime}}(\theta)-y_{k,h}^{s^{\prime}})\phi_{k,h}^{s^{\prime}}$, we have
$\lVert g_{k,h}(\widetilde{\theta}_{k,h})\rVert_{2}\leq\Big{\lVert}\sum_{s^{\prime}\in\mathcal{S}_{k,h}}p_{k,h}^{s^{\prime}}(\widetilde{\theta}_{k,h})\phi_{k,h}^{s^{\prime}}\Big{\rVert}_{2}+\Big{\lVert}\sum_{s^{\prime}\in\mathcal{S}_{k,h}}y_{k,h}^{s^{\prime}}\phi_{k,h}^{s^{\prime}}\Big{\rVert}_{2}\leq 2\max_{s^{\prime}\in\mathcal{S}_{k,h}}\big{\lVert}\phi_{k,h}^{s^{\prime}}\big{\rVert}_{2}\leq 2.$
Therefore, we have
$\sum_{k=1}^{K}\sum_{h=1}^{H}\sum_{s^{\prime}\in\mathcal{S}_{k,h}}p_{k,h}^{s^{\prime}}(\widetilde{\theta}_{k,h})\big{\lVert}\bar{\phi}_{k,h}^{s^{\prime}}-\widetilde{\phi}_{k,h}^{s^{\prime}}\big{\rVert}_{\mathcal{H}_{k,h}^{-1}}\leq$ $\frac{4\eta}{\sqrt{\lambda}}\sum_{k=1}^{K}\sum_{h=1}^{H}\sum_{s^{\prime}\in\mathcal{S}_{k,h}}p_{k,h}^{s^{\prime}}(\widetilde{\theta}_{k,h})\max_{s^{\prime\prime}\in\mathcal{S}_{k,h}}\big{\lVert}\phi_{k,h}^{s^{\prime\prime}}\big{\rVert}_{\mathcal{H}_{k,h}^{-1}}^{2}$
[…]
where the last inequality holds by Lemma 16.
For the term $(d)$, by similar analysis, we have
$(p_{k,h}^{s^{\prime}}(\widetilde{\theta}_{k,h})-p_{k,h}^{s^{\prime}}(\widetilde{\theta}_{k+1,h}))\big{\lVert}\widetilde{\phi}_{k,h}^{s^{\prime}}\big{\rVert}_{\mathcal{H}_{k,h}^{-1}}$
=\ = $\nabla p_{k,h}^{s^{\prime}}(\xi_{k,h})^{\top}(\widetilde{\theta}_{k,h}-\widetilde{\theta}_{k+1,h})\big{\lVert}\widetilde{\phi}_{k,h}^{s^{\prime}}\big{\rVert}_{\mathcal{H}_{k,h}^{-1}}$
=\ = $\Big{(}p_{k,h}^{s^{\prime}}(\xi_{k,h})\phi_{k,h}^{s^{\prime}}-p_{k,h}^{s^{\prime}}(\xi_{k,h})\sum_{s^{\prime\prime}\in\mathcal{S}_{k,h}}p_{k,h}^{s^{\prime\prime}}(\xi_{k,h})\phi_{k,h}^{s^{\prime\prime}}\Big{)}^{\top}(\widetilde{\theta}_{k+1,h}-\widetilde{\theta}_{k,h})\big{\lVert}\widetilde{\phi}_{k,h}^{s^{\prime}}\big{\rVert}_{\mathcal{H}_{k,h}^{-1}}$
[…]
where the third inequality holds by the AM-GM inequality. Thus, we have
$\sum_{k=1}^{K}\sum_{h=1}^{H}\sum_{s^{\prime}\in\mathcal{S}_{k,h}}(p_{k,h}^{s^{\prime}}(\widetilde{\theta}_{k,h})-p_{k,h}^{s^{\prime}}(\widetilde{\theta}_{k+1,h}))\big{\lVert}\widetilde{\phi}_{k,h}^{s^{\prime}}\big{\rVert}_{\mathcal{H}_{k,h}^{-1}}$
[…]
where the last inequality holds by Lemma 16.
Finally, we bound the term $(e)$ as follows.
$\sum_{k=1}^{K}\sum_{s^{\prime}\in\mathcal{S}_{k,h}}p_{k,h}^{s^{\prime}}(\widetilde{\theta}_{k+1,h})\big{\lVert}\widetilde{\phi}_{k,h}^{s^{\prime}}\big{\rVert}_{\mathcal{H}_{k,h}^{-1}}\leq$ $\sqrt{\sum_{k=1}^{K}\sum_{s^{\prime}\in\mathcal{S}_{k,h}}p_{k,h}^{s^{\prime}}(\widetilde{\theta}_{k+1,h})}\sqrt{\sum_{k=1}^{K}\sum_{s^{\prime}\in\mathcal{S}_{k,h}}p_{k,h}^{s^{\prime}}(\widetilde{\theta}_{k+1,h})\big{\lVert}\widetilde{\phi}_{k,h}^{s^{\prime}}\big{\rVert}_{\mathcal{H}_{k,h}^{-1}}^{2}}$
[…]
where the last inequality holds by Lemma 16.
Thus, combining (32), (33), and (34), we have
$\sum_{k=1}^{K}\sum_{h=1}^{H}\epsilon_{k,h}^{\mathtt{fst}}$ $=H\widetilde{\beta}_{K}\sum_{k=1}^{K}\sum_{h=1}^{H}\sum_{s^{\prime}\in\mathcal{S}_{k,h}}p_{k,h}^{s^{\prime}}(\widetilde{\theta}_{k,h})\Big{\lVert}\phi_{k,h}^{s^{\prime}}-\sum_{s^{\prime\prime}\in\mathcal{S}_{k,h}}p_{k,h}^{s^{\prime\prime}}(\widetilde{\theta}_{k,h})\phi_{k,h}^{s^{\prime\prime}}\Big{\rVert}_{\mathcal{H}_{k,h}^{-1}}$
$\leq H^{2}\widetilde{\beta}_{K}\left(\sqrt{2dK\log\left(1+\frac{K}{d\lambda}\right)}+\frac{16\eta}{\kappa\sqrt{\lambda}}d\log(1+\frac{K}{d\lambda})\right)$ (35)
Bounding $\epsilon_{k,h}^{\mathtt{snd}}$.
For the second-order term, by Lemma 16, we have
$\sum_{k=1}^{K}\sum_{h=1}^{H}\epsilon_{k,h}^{\mathtt{snd}}=\frac{5}{2}H\widetilde{\beta}_{k}^{2}\sum_{k=1}^{K}\sum_{h=1}^{H}\max_{s^{\prime}\in\mathcal{S}_{k,h}}\lVert\phi_{k,h}^{s^{\prime}}\rVert_{\mathcal{H}_{k,h}^{-1}}^{2}\leq\frac{5}{\kappa}H^{2}\widetilde{\beta}_{K}^{2}d\log\left(1+\frac{K}{d\lambda}\right).$ (36)
Combining (35) and (36), we have
$\operatorname*{Reg}(K)$ $\leq 2\sum_{k=1}^{K}\sum_{h=1}^{H}(\epsilon_{k,h}^{\mathtt{fst}}+\epsilon_{k,h}^{\mathtt{snd}})+H\sqrt{2KH\log(2/\delta)}$
$\leq H^{2}\widetilde{\beta}_{K}\left(\sqrt{2dK\log\left(1+\frac{K}{d\lambda}\right)}+\frac{16\eta}{\kappa\sqrt{\lambda}}d\log(1+\frac{K}{d\lambda})\right)+\frac{5}{\kappa}H^{2}\widetilde{\beta}_{K}^{2}d\log\left(1+\frac{K}{d\lambda}\right)$
$\hskip 28.45274pt+H\sqrt{2KH\log\Big{(}\frac{2}{\delta}\Big{)}}$
$\leq\widetilde{\mathcal{O}}\big{(}dH^{2}\sqrt{K}+d^{2}H^{2}\kappa^{-1}\big{)}$
This finishes the proof. ∎
Appendix I Proof of Theorem 4
Proof.
Our proof is similar to the proof of the lower bound for linear mixture MDPs in Zhao et al. [2023]. We prove this lemma by translating this problem to learning $H/2$ logistic bandit problems.
Specifically, we use each three layers to construct a block. Note that the third layer of block $i$ is also the first layer of block $i+1$ and hence there are total $H/2$ blocks. In each block, both the first and third layers of this block only have one state, and the second layer has two states. Here we take block $i$ as an example. The first two layers of this block are associated with transition probability $\mathbb{P}_{i,1}$ and $\mathbb{P}_{i,2}$. Denote by $s_{i,1}$ the only state in the first layer of this block. In the second layer of the block $i$, we assume there exist two states $s_{i,2}^{*}$ and $s_{i,2}$. Let $s_{i,3}$ be the only state in the third layer of this block. Further, for any $a\in\mathcal{A}$, let $\phi(s_{i,1},a,s_{i,2})=0$. The probability of transferring to state $s_{i,2}^{*}$ when executing action $a$ at state $s_{i,1}$ is $\rho_{a}$, in particular,
$\mathbb{P}_{i,1}(s_{i,2}^{*}\mid s_{i,1},a)$ $=\frac{\exp(\phi(s_{i,2}^{*}\>|\>s_{i,1},a)^{\top}\theta_{i,1}^{*})}{1+\exp(\phi(s_{i,2}^{*}\>|\>s_{i,1},a)^{\top}\theta_{i,1}^{*})}=\rho_{a},\quad\mathbb{P}_{i,1}(s_{i,2}\mid s_{i,1},a)=1-\rho_{a}.$
For the second layer, $\forall s=s_{i,2}^{*},s_{i,2}$, and $a\in\mathcal{A}$, $\mathbb{P}_{i,2}(s_{i,3}\mid s,a)=1$. The reward satisfies $r_{k}(s_{i,1},a)=0$ for the first layer and $r_{k}(s_{i,2}^{*},a)=1$, $r_{k}(s_{i,2},a)=0$ for the second layer for all $a\in\mathcal{A}$.
Consider the logistic bandit problem where a learner selects action $x\in\mathbb{R}^{d}$ and receives a reward $r_{k}$ sampled from Bernoulli distributed with mean $\mu(x^{\top}\theta^{*})=(1+\exp(-x^{\top}\theta^{*})^{-1})$. Thus, learning in this block of MDP can be regarded as learning a $d$-dimensional logistic bandit problem with $A$ arms, where the arm set is $\phi(s_{i,2}^{*}\>|\>s_{i,1},a)$ and the expected reward of each arm is $\rho_{a}$. It is also clear that the optimal policy at state $s_{i,1}$ is to select action $a_{i,1}^{*}=\operatorname*{arg\,max}_{a\in\mathcal{A}}\phi(s_{i,2}^{*}\>|\>s_{i,1},a)$. Thus, learning this MDP is equal to learning $H/2$ logistic bandit problems. This finishes the proof. ∎
Appendix J Supporting Lemmas
In this part, we provide several supporting lemmas used in the proofs of the main results.
Lemma 15 (Abbasi-Yadkori et al. [2011, Theorem 1]).
Let $\left\{\mathcal{F}_{t}\right\}_{t=0}^{\infty}$ be a filtration. Let $\left\{\eta_{t}\right\}_{t=1}^{\infty}$ be a real-valued stochastic process such that $\eta_{t}$ is $\mathcal{F}_{t}$-measurable and $\eta_{t}$ is conditionally zero-mean $R$-sub-Gaussian for some $R\geq 0$ i.e. $\forall\lambda\in\mathbb{R}$, $\mathbb{E}\left[e^{\lambda\eta_{t}}\mid F_{t-1}\right]\leq\exp\left({\lambda^{2}R^{2}}/{2}\right)$. Let $\left\{X_{t}\right\}_{t=1}^{\infty}$ be an $\mathbb{R}^{d}$-valued stochastic process such that $X_{t}$ is $\mathcal{F}_{t-1}$-measurable. Assume that $V\in\mathbb{R}^{d\times d}$ is a positive definite matrix. For any $t\geq 0$, define $V_{t}=V+\sum_{s=1}^{t-1}X_{s}X_{s}^{\top},S_{t}=\sum_{s=1}^{t-1}\eta_{s}X_{s}$. Then, for any $\delta>0$, with probability at least $1-\delta$, for all $t\geq 0$,
$\left\|S_{t}\right\|_{V_{t}^{-1}}^{2}\leq 2R^{2}\log\left(\frac{\det\left(V_{t}\right)^{1/2}\det(V)^{-1/2}}{\delta}\right).$
Lemma 16 (Abbasi-Yadkori et al. [2011, Lemma 10]).
Suppose $x_{1},\ldots,x_{t}\in\mathbb{R}^{d}$ and for any $1\leq s\leq t$, $\lVert x_{s}\rVert_{2}\leq L$. Let $V_{t}=\lambda I_{d}+\sum_{s=1}^{t}x_{s}x_{s}^{\top}$ for $\lambda\geq 0$. Then, we have $\det(V_{t})\leq(\lambda+{tL^{2}}/{d})^{d}$.
Lemma 17 (Orabona [2019, Lemma 6.9]).
Let $Z$ be a positive define matrix and $\mathcal{X}$ be a convex set, define $\widetilde{\mathbf{x}}=\operatorname*{arg\,min}_{\mathbf{x}\in\mathcal{X}}\langle\mathbf{g},\mathbf{x}\rangle+\frac{1}{2\eta}\lVert\mathbf{x}-\widehat{\mathbf{x}}\rVert_{Z}^{2}$. Then, for all $\mathbf{x}^{*}\in\mathcal{X}$, we have
$\lVert\widetilde{\mathbf{x}}-\widehat{\mathbf{x}}\rVert_{Z}\leq 2\eta\lVert\mathbf{g}\rVert_{Z^{-1}}\quad\mbox{and}\quad\langle\mathbf{g},\widehat{\mathbf{x}}-\mathbf{x}^{*}\rangle\leq\frac{\eta}{2}\lVert\mathbf{g}\rVert_{Z^{-1}}^{2}+\frac{1}{2\eta}\lVert\widehat{\mathbf{x}}-\mathbf{x}^{*}\rVert_{Z}^{2}-\frac{1}{2\eta}\lVert\widetilde{\mathbf{x}}-\mathbf{x}^{*}\rVert_{Z}^{2}.$
Lemma 18 (Lee and Oh [2024, Lemma D.3]).
Define $Q:\mathbb{R}^{K}\rightarrow\mathbb{R}$, such that for any $\mathbf{u}=\left(u_{1},\ldots,u_{K}\right)\in\mathbb{R}^{K},Q(\mathbf{u})=$ $\sum_{i=1}^{K}\frac{\exp\left(u_{i}\right)}{v+\sum_{k=1}^{K}\exp\left(u_{k}\right)}$. Let $p_{i}(\mathbf{u})=\frac{\exp\left(u_{i}\right)}{v+\sum_{k=1}^{K}\exp\left(u_{k}\right)}$. Then, for all $i\in[K]$,
$\left|\frac{\partial^{2}Q}{\partial i\partial j}\right|\leqslant\begin{cases}3p_{i}(\mathbf{u})&\text{ if }i=j,\\ 2p_{i}(\mathbf{u})p_{j}(\mathbf{u})&\text{ if }i\neq j.\end{cases}$